### v1

# EF-3DGS: Event-Aided Free-Trajectory 3D Gaussian Splatting

Abstract
Scene reconstruction from casually captured videos has wide applications in real-world scenarios. With recent advancements in differentiable rendering techniques, several methods have attempted to simultaneously optimize scene representations (NeRF or 3DGS) and camera poses. Despite recent progress, existing methods relying on traditional camera input tend to fail in high-speed (or equivalently low-frame-rate) scenarios due to insufficient observations and large pixel displacements between adjacent frames. Event cameras, inspired by biological vision, record pixel-wise intensity changes asynchronously with high temporal resolution and low latency, providing valuable scene and motion information in blind inter-frame intervals. In this paper, we introduce the event camera to aid scene construction from a casually captured video for the first time, and propose Event-Aided Free-Trajectory 3DGS, called EF-3DGS, which seamlessly integrates the advantages of event cameras into 3DGS through three key components. First, we leverage the Event Generation Model (EGM) to fuse events and frames, supervising the rendered views observed by the event stream. Second, we adopt the Contrast Maximization (CMax) framework in a piece-wise manner to extract motion information by maximizing the contrast of the Image of Warped Events (IWE), thereby calibrating the estimated poses. Besides, based on the Linear Event Generation Model (LEGM), the brightness information encoded in the IWE is also utilized to constrain the 3DGS in the gradient domain. Third, to mitigate the absence of color information of events, we introduce photometric bundle adjustment (PBA) to ensure view consistency across events and frames. Additionally, we propose a Fixed-GS training strategy that separates the optimization of scene structure and color, effectively addressing color distortions caused by the lack of color information in events. We evaluate our method on the public Tanks and Temples benchmark and a newly collected real-world dataset, RealEv-DAVIS. Our method achieves up to 2dB higher PSNR and 40% lower Absolute Trajectory Error (ATE) compared to state-of-the-art methods under challenging high-speed scenarios. Our project page is https://lbh666.github.io/ef-3dgs/.
Index Terms:
Event Camera, Novel View Synthesis, 3D Gaussian Splatting, Neural Rendering
Figure 1: Free-trajectory 3DGS under high speed. The top row shows the overall paradigm. The colored dots (red for positive and blue for negative events) in the top row represent the event data. Given free-trajectory video and corresponding event stream under the high-speed scenario, we aim to jointly estimate poses and the 3DGS. Our method surpasses current state-of-the-art methods in terms of both rendered results (middle row) and pose estimation (bottom row).
I Introduction
In recent years, Neural Radiance Fields (NeRF) [1, 2, 3] and 3D Gaussian splatting (3DGS) [4, 5] have made significant progress in novel view synthesis tasks. Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering. While subsequent methods [2, 3, 6, 7] have made great strides in rendering quality and speed, how to reconstruct the scene from a video captured with free camera trajectories remains an open problem. It has great application potential in the real world, such as in VR/AR, video stabilization, and mapping. To tackle this challenging task, several efforts have been made.
Free-trajectory videos often cover a vast field of view, so a key challenge is how to preserve the detailed textures of the foreground while representing the overall large-scale scene. Previous methods [8, 9] adopt different spatial transformation and allocation approaches based on neighboring frames’ viewing angles and range to achieve efficient and high-quality scene reconstruction. Another challenge is that accurate pose estimation is often difficult to obtain in free-trajectory scenarios. Previous methods [9, 10] have incorporated priors, such as the depth [11, 12, 13] and the optical flow [14], into their loss functions to improve pose estimation or modify their optimization strategies. While these methods above can render photo-realistic images in typical free-trajectory scenarios, the rendering quality degrades significantly in high-speed scenarios (or equivalent low-frame-rate scenarios). This type of scenario has essential applications in the real world, such as autonomous driving or FPV exploration. Besides, the significant motion between consecutive frames poses challenges for feature matching and further abrupt pose estimation (as shown in the third row of Fig.1). This makes it even more crucial for the algorithm to robustly estimate the poses in such scenarios.
The performance degradation of prior methods can be attributed to two primary factors. Firstly, the limited number of camera observations leads to an under-constrained scene reconstruction problem. This can cause the scene representation to converge to a trivial solution [15, 16, 17], where the model overfits to the training views without capturing the correct underlying geometry structure. Consequently, the synthesized images from novel viewpoints exhibit visual artifacts. Secondly, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which was leveraged by previous techniques such as LocalRF [9] and NoPe-NeRF [10]. This significant violation greatly exacerbates the ill-posedness of the joint optimization of scene and camera poses, further deteriorating the quality of the synthesized novel views.
Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur [18, 19, 20, 21, 22, 23]. The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames. Moreover, in scenarios with stable scene illumination, the event data naturally encodes the motion information of the scene [24, 25, 26], containing rich motion cues. These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios [27]. However, seamlessly integrating the aforementioned benefits of event cameras is nontrivial. First, NeRF or 3DGS renders absolute pixel brightness, which aligns with image data. In contrast, event stream records sparse scene brightness changes, lacking the ability to perceive absolute brightness, and most event cameras lack color information. Second, the motion cues in event cameras is encoded in the spatio-temporal correlations of the event stream. In contrast, previous methods typically sample and optimize frames and pixels independently, without explicitly considering these temporal correlations.
In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage the high temporal resolution property of event cameras. Specifically, based on the Event Generation Model (GEM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream. This allows us to effectively utilize the brightness cues encoded in the event streams between adjacent frames, providing rich supervisory signals to alleviate the problem of insufficient sparse viewpoints. Regarding the pose estimation challenge, we introduce the CMax [28] framework to exploit the spatio-temporal correlations of events. We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive time. We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns. Furthermore, through the Linear Event Generation Model (LEGM) [29, 30], we establish a connection between motion and brightness changes. This allows us to constrain the 3DGS in the gradient domain using the IWE. Finally, as event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this. PBA recovers color by optimizing reprojection errors onto RGB frames, while Fixed-GS enables separate optimization of scene structure and color.
Our main contributions are summarized as follows:
  • We introduce event cameras into the task of free-trajectory scene reconstruction for the first time. Its advantage of high temporal resolution and low latency showcases the potential of event data for scene reconstruction tasks in challenging scenarios.
  • We derive our method from the underlying imaging principles of event cameras and design the corresponding loss functions that mine the motion and brightness information encoded in event data and seamlessly integrate them into the 3DGS optimization.
  • Experiments on both public benchmarks and real-world datasets demonstrate that our method significantly outperforms existing state-of-the-art approaches in terms of both rendering quality and trajectory estimation accuracy. Our method achieves up to 2dB higher PSNR and 40% lower Absolute Trajectory Error compared to state-of-the-art methods in high-speed scenarios.
II RELATED WORKS
II-A Large Scale Scene Reconstruction
The reconstruction of large scenes has emerged as a significant area of research over the past decade. Traditional approaches [31, 32] rely on Structure from Motion techniques to recover camera poses and sparse 3D point clouds. The introduction of NeRF [1] marked a paradigm shift in scene representation and reconstruction. Several works [33, 34, 2] have attempted to extend NeRF to unbounded scenes. F2-NeRF [8] adopts perspective warping based on the local viewing range to handle free-viewpoint unbounded scenes. To extend unbounded scenes to large-scale scenes, Mega-NeRF [35] and Block-NeRF [36] adopt a divide-and-conquer strategy. LocalRF [9] allocates new radiance fields dynamically according to the distance between the current view and the previous radiance field. Recently, 3D Gaussian Splatting (3DGS) has been extended to large-scale scene reconstruction due to its efficient rendering capabilities and explicit 3D representation. CityGaussian [37] and VastGaussian [38] adopt spatial allocation strategies similar to Mega-NeRF [35]. CityGaussian [37] dynamically selects poses for optimization based on visibility criteria, while VastGaussian [38] employs a visibility-driven progressive approach to dynamically select both training poses and Gaussian points for optimization. In contrast to these methods, which primarily focus on scalability and often require precomputed camera poses from COLMAP [39], we focus on the challenges of reconstructing scene from casually captured free-trajectory videos without relying on precomputed poses.
II-B Joint Pose and Scene Optimization
The research community has recently focused on developing methods [17, 9, 40, 41, 10, 42, 5] that can be optimized without requiring precomputed camera poses. A line of work has focused on improving the stability of the optimization process. GARF [41] and BARF [42] both find that the high-frequency position encoding is prone to local minima and try to improve it. For example, GARF [41] proposes using Gaussian activation to replace the sinusoidal position encoding. Another line of work has investigated incorporating additional constraints to make the problem more tractable. SPARF [17] incorporates cross-view feature matching constraints and depth consistency regularization. More recent approaches [10, 9, 5] leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation, to obtain additional priors and better address the pose-free NeRF reconstruction problem. Exploiting 3DGS’s explicit representation, CF-3DGS [5] directly back-projects Gaussian points using depth maps, enhancing the utilization of depth priors. While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed ( or equivalently sparse-view) scenarios or rely on a good pose initialization. Our approach addresses these issues by leveraging motion and brightness cues from event streams.
II-C Event-Based Novel view Synthesis
Recent works have explored the integration of event cameras into the NeRF framework. Early approaches, such as E-NeRF [43] and EventNeRF [44], utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes calculated by the accumulated events. Building upon this, Robust e-NeRF [45] incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise. Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods. For instance, E2NeRF [46] and EvDeblurNeRF [47] leverage the Event Double Integral (EDI) [48] model to address the deblurring problem, while DE-NeRF [49] and EvDNeRF [50] leverage the high temporal resolution property of event camera to capture fast-moving elements in dynamic scene. A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we jointly optimize for both the camera poses and the scene representation. Furthermore, while previous works have been limited to simulated environments or simple indoor scenes, we evaluate our approach in large-scale outdoor scenarios.
III Preliminary
3DGS [4] parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information. Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$. The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$,
$G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ (1)
To preserve the valid positive semi-definite property during optimization, the $\mathbf{\Sigma}$ is derived by the scale factor $s$ and the rotation factor $R$,
$\mathbf{\Sigma}=RSS^{T}R^{T}.$ (2)
Before rendering, the Gaussian is first projected to the 2D image plane:
$\mathbf{\Sigma}^{2D}=JW\mathbf{\Sigma}W^{T}J^{T},$ (3)
where J is the Jacobian of the affine approximation of the projective transformation, and W is the viewing transformation. During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending. The color of each pixel is calculated via blending N ordered overlapping points:
$C(\textbf{r})=\sum_{i=1}^{N}c_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}),$ (4)
where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel. With the forward rendering procedure, we can optimize 3DGS by minimizing the photometric loss between observation and rendered pixels:
$\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ (5)
where the first term is a standard $\mathcal{L}_{1}$ loss between the target and rendered images, $\mathcal{L}_{D-SSIM}$ is a differentiable version of structural similarity index and $\lambda$ is balancing weight which is set to 0.2 following [4]. As this optimization is highly sensitive to local minima, 3DGS additionally employs periodic adaptive densification and pruning of splats. By integrating depth $t_{i}$ of each Gaussian in Equation 4 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
$\hat{D}(\mathbf{r})=\sum_{i=1}^{N}t_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}).$ (6)
Figure 2: Method Overview. Our method takes video frames $\{I_{i}\}$ and event stream $\varepsilon=\{\mathbf{e}_{k}\}$ as input. During training, we random sample $t\in\{t_{i,j}\}$, leveraging the events within current subinterval and most recent frame to establish the proposed three main constrains,$\mathcal{L}_{EGM}$, $\mathcal{L}_{LEGM}$ and $\mathcal{L}_{PBA}$. The colored dots (red for positive and blue for negative events) represent the event data. The details of $\mathcal{L}_{LEGM}$ and CMax framework are illustrated in Fig. 3
IV Method
In this section, we introduce our method. The overall framework is shown in Fig. 2. Given a video of a free-trajectory $\{I_{i}\}$ captured at time $\{t_{i}\}$ and an event stream $\varepsilon=\{\mathbf{e}_{k}\}$, our goal is to reconstruct the 3DGS of the scene and the corresponding camera trajectory$\{T_{i}\}$. The photometric loss of the original 3DGS follows an analysis-by-synthesis process. Similarly, in order to incorporate the events into the 3DGS optimization process, we delve into the imaging principles of event cameras. Revolving around the Event Generative Model (EGM) and Linearized Event Generative Model (LEGM), we establish constraints from the connection between motion and brightness changes. Further, through photometric bundle adjustment (PBA) and Fix-GS training strategy, we mitigate the absence of color information of events and further strengthen the consistency of cross-view rendering.
IV-A EGM Driven Optimization
The Event Generation Model (EGM) describes the base working principle of event cameras, which capture pixel-wise brightness changes asynchronously. When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,
$\Delta L(\mathbf{u}_{k},t_{k})\doteq L(\mathbf{u}_{k},t_{k})-L(\mathbf{u}_{k},t_{k}-\delta t)=p_{k}C,$ (7)
where $L\doteq log(I)$ is the logarithm of intensity, $p_{k}\in\{-1,+1\}$ indicates the polarity of brightness changes, and $t_{k}$ is the triggered timestamp. By leveraging the high temporal resolution of event cameras, we can effectively constrain the 3DGS optimization process in the temporal domain.
To incorporate the event data into the optimization process, we first divide the time interval between two adjacent video frames $I_{i}$ and $I_{i+1}$ into N smaller subintervals $\varepsilon_{i,j}=\{\mathbf{e}_{k}|t_{i,j}\leq t_{k}\leq t_{i,j+1},\Delta t=\frac{t_{i+1}-t_{i}}{N},t_{i,j}=t_{i}+j\cdot\Delta t\}$. This allows us to form event frames at a higher temporal resolution than the original video frames:
$E_{i,j}=\sum_{e_{k}\in\varepsilon_{i,j}}p_{k}.$ (8)
We then reconstruct the latent intensity image $I_{t}$ at any intermediate time $t\in\{t_{i,j}\}$ between two video frames $I_{i}$ and $I_{i+1}$ using the accumulated events and the most recent frame:
$I_{t}=I_{i,j}=\begin{cases}I_{i,0}\cdot\exp(\sum_{n=0}^{j-1}E_{i,n}\cdot C)&\text{if }j>0\\ I_{i,0}&\text{if }j=0\end{cases}.$ (9)
Then, the event-based rendering loss is introduced to enforce the consistency between rendered and latent intensity images:
$\mathcal{L}_{EGM}=(1-\lambda)\mathcal{L}_{1}(\hat{I}_{t},I_{t})+\lambda\mathcal{L}_{D-SSIM}(\hat{I}_{t},I_{t}).$ (10)
This loss function ensures that the 3DGS optimization process is constrained not only by the sparse video frames but also by the dense event data, leading to improved reconstruction quality. It is worth noting that when $j=0$, the latent intensity image $I_{t}$ is equal to the first most recent video frame $I_{i,0}$ and $\mathcal{L}_{EGM}$ reduces to the standard image rendering loss.
IV-B Unified CMax and LEGM Optimization
While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by the events, it does not explicitly exploit the motion information encoded in the event stream. To further improve the accuracy of camera poses and the reconstructed 3DGS, we introduce the Contrast Maximization (CMax) [28, 51, 52] framework and the Linear Event Generation Model (LEGM) [29, 30, 53] in this section. These models complement the EGM-driven optimization by extracting motion cues and gradient information from the event stream, providing additional constraints for the 3DGS optimization process.
Under the assumption of constant scene illumination, events are triggered solely by the motion of scene edges. The edges in the scene form continuous trajectories in the $(x,y,t)$ space. Intuitively, by back-projecting the events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE). Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the currently estimated motion. This insight motivates us to derive a motion field by leveraging the rendered depth from 3DGS using eq. 6 and the relative camera motion between neighboring timestamps. By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the relative poses between adjacent frames.
Instead of warping each event as in the original CMax framework, we adopt a piece-wise warping approach for efficiency. Specifically, we first accumulate the events within each sub-interval to obtain event frames $E_{i,j}$ as in eq. 8 . Then, for randomly sampled timestamps $t_{ref}=t_{i,j}$, we warp the event frames in the neighboring $2r$ sub-intervals to $t_{ref}$:
$E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),$ (11)
where $m\in[-r,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow which can be derived from the rendered depth $\hat{D}$ in  eq. 6 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
$F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),$ (12)
where $T_{i,j\rightarrow j-m}=T_{i,j-m}T_{i,j}^{-1}$, $\Pi$ projects a 3D point to image coordinates and $\Pi^{-1}$ unprojects a pixel coordinate and depth into a 3D point. Then the image of piece-wise warped events (IPWE) at timestamp $t_{i,j}$ is computed by averaging the warped event frames:
$\textrm{IPWE}_{i,j}=\frac{1}{2r+1}\sum_{m=j-r}^{j+r}E_{i,m\rightarrow j}\approx\frac{1}{C}\Delta L.$ (13)
Following the Cmax framework, we maximize the variance of the IPWE, which is equivalent to minimize its opposite:
$\mathcal{L}_{cm}=-\textrm{Var}(\textrm{IPWE}_{i,j}).$ (14)
Maximizing the contrast encourages the motion field to align with the actual motion of edges, leading to sharper IPWEs. This objective helps to improve the accuracy of the estimated camera poses and the geometric consistency of the 3DGS reconstruction. Furthermore, based on the Linear Event Generation Model (LEGM) [29, 30], the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of a image):
$\Delta L(\mathbf{u})=-\nabla L\cdot\dot{\mathbf{u}}\approx L(\mathbf{u})-L(\mathbf{u}+\dot{\mathbf{u}}).$ (15)
It is noteworthy that the IPWE also encodes brightness change information. Combining  eq. 13 and  eq. 15, we establish a connection between the IPWE and the brightness changes of the rendered images:
$C\cdot\textrm{IPWE}_{i,j}=\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}).$ (16)
Based on this relationship, we formulate an additional gradient-based loss to constrain the 3DGS optimization:
$\mathcal{L}_{grad}=||C\cdot\textrm{IPWE}_{i,j}-(\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}))||^{2},$ (17)
where $\hat{L}$ is the logarithm of synthesised image $\hat{I}_{t}$. Finally, the full LEGM loss is defined as:
$\mathcal{L}_{LEGM}=\lambda_{cm}\mathcal{L}_{cm}+\lambda_{grad}\mathcal{L}_{grad},$ (18)
where $\lambda_{cm}$ and $\lambda_{grad}$ are the balancing weight coefficients for $\mathcal{L}_{cm}$ and $\mathcal{L}_{grad}$ respectively.
IV-C Photometric Bundle Adjustment
The aforementioned event-based constrains, $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$, leverage the brightness change and motion information encoded in the event data to constrain 3DGS. However, as event cameras only record brightness changes and lack color perception, directly applying them to 3DGS optimization may lead to inconsistent color rendering. To ensure cross-view consistency of the 3DGS rendering, we introduce the Photometric Bundle Adjustment (PBA) term.
Specifically, for a randomly sampled timestamp $t\in\{t_{i,j}\}$, we establish the following photometric reprojection error:
$\mathcal{L}_{PBA}=\sum_{\mathbf{u}\in\mathcal{P}}\sum_{I_{s}\in\mathcal{F}}||I_{s}(\mathbf{u}^{\prime})-\hat{I}(\mathbf{u})||^{2},$ (19)
where $\mathbf{u}^{\prime}=\Pi(T_{i,j-r\rightarrow j}\Pi^{-1}(x,y,\hat{D}(u)))$ represent the coordinate on target view projected from the pixel $\mathbf{u}$ of source view $I_{s}$, $\mathcal{P}$ denotes the pixel samples of current frame, and $\mathcal{F}$ is the candidates of target video frames. We select $\mathcal{F}$ to be the adjacent video frames in consideration of computation costs.
By minimizing $\mathcal{L}_{PBA}$ over sampled pixels in the event frames and neighboring video frames, we encourage the 3DGS model to produce geometrically and photometrically consistent renderings across events and video frames.
Figure 3: The illustration of unified CMax and LEGM optimization. We warp nearby event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of piece-wise warped events (IPWE). The byproduct IWPE is utilized to establish additional constrain on 3DGS in gradient domain.
IV-D Fixed-GS Training Strategy
As demonstrated in Fig. 10, the $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues. The absence of color information in event data poses a significant challenge for scene reconstruction in previous NeRF-based methods. NeRF models employ implicit representations [54], utilizing MLPs to store scene color and density information in a coupled manner. Consequently, color distortions can affect scene density estimations and vice versa. In contrast, 3DGS explicitly stores various attributes independently within Gaussian points. This characteristic allows us to address color distortion issues specifically while maintaining other attributes.
To leverage this advantage, we propose a two-stage Fixed-GS scene optimization strategy. In the first stage, optimization is performed across all timestamps:
$G_{\theta}^{*}=\underset{\alpha,r,s,f}{\operatorname{argmin}}~{}\mathcal{L}_{event},~{}t\in\{t_{i,j}\},$ (20)
where $\alpha,r,s,f$ is the opacity, rotation, scale factor and spherical harmonics of Gaussians, and t is the sampled timestamp during training. This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions. The second stage focuses on recovering accurate color information. During this phase, optimization is conducted exclusively on video frames. Capitalizing on the explicit representation of Gaussians, we optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed.
[…]
The ratio between the first and second stages is empirically set to 4:1. This approach allows us to effectively address the color distortion problem while preserving the structural and brightness information obtained from the event data.
IV-E Overall Training Pipeline
Assembling all loss terms, we get the overall loss function for events:
$\mathcal{L}_{event}=\mathcal{L}_{EGM}+\mathcal{L}_{LEGM}+\lambda_{PBA}\mathcal{L}_{PBA},$ (22)
where $\lambda_{PBA}$ are the weighting factor. It is noteworthy that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space. We borrow the idea of dynamic scene allocation strategies from LocalRF [9]. Our overall training pipeline builds upon the pipeline of CF-3DGS [5] while introducing novel components to integrate event camera data for free-trajectory scene reconstruction. Further implementation details are provided in the Supplementary Material.
V Experiments
TABLE I: Quantitative Evaluations on Tanks and Temples. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are highlighted in bold.
Methods Pose-Free Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
[…]
F2-NeRF $\times$ F 23.55 0.75 0.34 22.97 0.72 0.36 22.25 0.69 0.40 21.64 0.68 0.44 20.63 0.64 0.51
[…]
EvDeblurNeRF $\times$ E+F 22.43 0.71 0.38 21.23 0.69 0.42 20.09 0.65 0.49 17.52 0.62 0.55 15.19 0.55 0.60
ENeRF $\times$ E+F 23.62 0.73 0.37 22.84 0.70 0.38 21.85 0.69 0.41 20.52 0.66 0.46 18.09 0.60 0.52
[…]
TABLE II: Pose accuracy on Tanks and Temples. Note that we use COLMAP poses in Tanks and Temples as the “ground truth”. The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100. Those methods which require precomputed poses are excluded.
Methods Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
[…]
Nope-NeRF F 0.1141 0.7563 2.8382 0.1604 1.0542 2.7653 0.2220 1.3694 2.7857 0.3131 1.8965 2.8412 0.6216 3.913 2.6592
LocalRF F 0.0806 0.9282 0.5630 0.0867 0.9683 0.6085 0.0911 0.9800 0.6501 0.0957 1.0428 0.6802 0.1421 1.4725 1.0006
CF-3DGS F 0.0594 0.6981 0.4212 0.0637 0.7128 0.4628 0.0712 0.7531 0.5189 0.0859 0.8074 0.6918 0.1057 0.9768 0.8972
Ev-Baseline E+F 0.0461 0.5972 0.3419 0.0490 0.6269 0.3766 0.0538 0.6728 0.4261 0.0591 0.7094 0.4860 0.0657 0.7597 0.5534
EF-3DGS(Ours) E+F 0.0391 0.5427 0.2885 0.0407 0.5521 0.3064 0.0426 0.5796 0.3271 0.0449 0.5953 0.3671 0.0487 0.6259 0.3753
Figure 4: Qualitative comparison for novel view synthesis on Tanks and Temples. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 5: Pose Estimation Comparison on Tanks and Temples. We visualise the trajectory (3D plot) and $\textrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
V-A Dataset
V-A1 Tanks and Temples
We conduct comprehensive experiments on the Tanks and Temples dataset [55]. Similar to LocalRF [9], we adopt the same 9 scenes, covering large-scale indoor and outdoor scenes. For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance. Following LocalRF [9], we spatially downsample the videos to a resolution of 1080p. Increasing motion speed makes frames further apart, hence it is effectively equivalent to reducing the frame rate. We employ various temporal downsamplings to evaluate the robustness of different methods to varying camera speed scenes. In the 6 FPS setting, we retain one in every five frames as in LocalRF [9]. We perform corresponding frame downsampling for the other frame rate settings (4 FPS, 3 FPS, 2 FPS, and 1 FPS). To synthesize realistic event data, we first upsample the original videos by [56] and then apply the simulator V2E [57].
V-A2 RealEv-DAVIS
Due to the current lack of datasets featuring free-trajectory captures from event cameras, we introduce the RealEv-DAVIS dataset, encompassing four outdoor scenes. We use the DAVIS346 camera, which simultaneously records image frames and events at a resolution of 346x260 pixels. To simulate real-world free-trajectory scenarios, we capture handheld footage at 25 FPS, with each sequence lasting 40 seconds. For SLOW scenarios, we downsample the video clips by keeping every second frame. We further reduce the frame rate in FAST scenarios by retaining only one frame out of every five.
V-B Implementation details
Our implementation is primarily based on the PyTorch framework, and we follow the optimization parameters by the configuration outlined in the 3DGS [4] unless otherwise specified. All the parameters are optimized using Adam [58] with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We optimize the camera poses in the representation of quaternion rotation. The initial learning rate is set to $10^{-5}$ and gradually decay to $10^{-6}$ until convergence. The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5, respectively, based on subsequent experiment results. For the division of events between adjacent frames, we maintain a constant interval of $\frac{1}{6}$s for Tanks and Temples and $\frac{1}{25}$s for RealEv-DAVIS, setting the number of subinterval N accordingly. For example, in Tanks and Temples, N equals 2 for 3FPS and 6 for 1FPS. This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for following CMax warping. The intervals of neighboring warping $r$ in Cmax are set to 3. That is, we warp the three adjacent event frames to the current sampled timestamp.
V-C Experimental Setup
V-C1 Metrics
We evaluate all the methods from two aspects: novel view synthesis and pose estimation. For the novel view synthesis task, we report the standard metrics PSNR, SSIM [59], and LPIPS [60] to quantitatively assess the quality of the rendered images. To measure the accuracy of pose estimation, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [61, 62], as delineated in [10]. Given that $\text{RPE}_{r}$ and $\text{RPE}_{t}$ measure relative pose errors between adjacent frames, they are inherently influenced by the frame rate. To facilitate a fair comparison across different frame rate settings, we upsample the estimated poses to the same frame rate before evaluation. This enables a consistent comparison of relative pose estimation errors across varying frame rate configurations.
TABLE III: Quantitative Evaluations on RealEv-DAVIS. We select the top-performing methods from the previous evaluation. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are highlighted in bold.
Methods Pose-Free Input SLOW FAST
hall building corner outdoor hall building corner outdoor
[…]
EvDeblurNeRF $\times$ E+F 20.24 0.5762 19.35 0.5649 20.47 0.5757 22.36 0.7087 18.23 0.5043 16.41 0.4624 17.82 0.4959 19.44 0.6448
ENeRF $\times$ E+F 21.89 0.5933 21.94 0.5918 22.01 0.5918 23.91 0.7266 20.09 0.5261 18.99 0.4903 19.61 0.5173 20.00 0.6494
[…]
TABLE IV: Pose accuracy on RealEv-DAVIS. The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100. Those methods which require precomputed poses are excluded.
Methods Input SLOW FAST
hall building corner outdoor hall building corner outdoor
$\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$
LocalRF F 4.4017 2.5067 4.6186 1.9409 1.9296 2.5183 3.4307 1.3005 5.6061 3.2265 7.0611 2.8722 3.4767 3.8950 4.7229 1.8595
CF-3DGS F 2.7106 0.6035 3.0374 3.2761 1.9437 1.4954 2.2747 0.8053 3.5025 0.7639 4.6758 4.5256 3.4950 2.2855 3.0551 1.1198
Ev-Baseline E+F 0.7063 0.4386 3.4225 1.1717 2.1283 1.3361 0.8647 0.3455 0.8898 0.7734 5.0707 1.6997 3.6785 2.1552 1.1645 0.5001
EF-3DGS(Ours) E+F 0.5262 0.3251 2.8008 0.4633 1.8386 1.6168 0.4726 0.3699 0.5424 0.3464 3.5848 0.5589 2.5396 2.2157 0.5218 0.4462
V-C2 Baselines
For a fair comparison, we focus on two categories of methods: those solely based on frames and those leveraging both events and frames. For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF and F2-NeRF, although the latter requires pre-estimated poses. We also include pose-free methods like Nope-NeRF and the current state-of-the-art CF-3DGS. To address the out-of-memory (OOM) issues CF-3DGS faces with long video sequences, we implement a slight modification, segmenting long videos into subsequences for individual reconstruction. For event-based methods, consistent with EvDeblurNeRF [47], we consider approaches that fuse events and frames. This category includes ENeRF [43] and the SOTA deblurring method EvDeblurNeRF [47]. These hybrid approaches leverage the complementary strengths of both data modalities. To the best of our knowledge, no existing method has exploit event cameras to address free-trajectory problem. To facilitate a fair comparison, we construct an event-based baseline method built upon CF-3DGS [5]. It leverages the high temporal resolution property of event cameras by utilizing an event-based frame interpolation network Time Lens [63] to interpolate the original video. Then, the temporally upsampled video frames are subsequently fed into CF-3DGS.
V-D Experimental Results
To evaluate the results on synthesized novel views, we select every ten frames as a test image following LocalRF [9]. Since the camera poses are unknown in our setting, we need to estimate the poses of test views. As in NeRFmm [40] and iNeRF [64], we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
V-D1 Results on Tanks and Temples
Quantitatively, the experimental results in Table I and Table II demonstrate the superiority of event-based methods (EF-3DGS and Ev-Baseline) over frame-based baselines in both pose estimation accuracy and rendering quality. This highlights the potential of event cameras for free-trajectory scene reconstruction. Furthermore, our method surpasses Ev-Baseline across all evaluation metrics under varying frame rate settings, indicating that our proposed method can better effectively integrate property of the high temporal resolution and motion information encoded in the event streams into the scene reconstruction process.
Qualitatively, as shown in Fig. 4, our method generates more realistic and detailed renderings compared to the baseline. The synthesized views of our method exhibit sharper edges and better preservation of finer textures. In contrast, the baseline methods suffer from various artifacts, such as blurriness and texture inconsistencies. Moreover, the visualization of estimated camera trajectories in Fig. 5 indicates that our estimated trajectory is better aligned with the ground truth than other methods, and the $\textrm{RPE}_{r}$ is significantly smaller.
Figure 6: Qualitative comparison for novel view synthesis on RealEv-DAVIS. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 7: Pose Estimation Comparison on RealEv-DAVIS. We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
V-D2 Results on RealEv-DAVIS
In Table III and Table IV, we report results on the self-collected dataset, RealEv-DAVIS. We select the top-performing methods from the previous evaluation. EF-3DGS demonstrates robust generalization to real-world scenarios, despite challenges such as sensor noise absent in simulated environments. Notably, the performance gap widens in FAST scenarios, with PSNR improvements of nearly 2 dB over baselines, emphasizing the method’s efficacy in handling high-speed motion in real-world settings. Qualitative results in Fig. 6 and Fig. 7 showcase EF-3DGS’s superior preservation of fine-grained details and the method’s robustness to real-world trajectory.
V-D3 Performance under Varying Camera Speeds
As evidenced in Table I to Table IV, the performance of all methods generally declines as the frame rate decreases. However, our method exhibits remarkable resilience to the reduction in frame rate, maintaining its superiority over the other methods across all metrics. Regarding novel view synthesis quality, the performance gap between the proposed method and the other methods consistently widens as the frame rate decreases. For example, the rendering quality difference between our method and CF-3DGS [5], measured in PSNR, increases from 0.61dB at 6FPS to 3.43dB at 1FPS. This observation demonstrates the robustness of EF-3DGS in handling challenging high-speed scenarios, where the advantages of leveraging event data become more pronounced.
V-D4 Robustness to Pose Disturbance
To further validate the performance of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP. Specifically, following BARF [42], we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra. For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level. Then, each method is initialized with the noised poses, after which the optimization is performed. The results are illustrated in Fig. 11. Notably, ENeRF [43], which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases. This observation validates the critical importance of optimizing camera poses alongside the scene reconstruction. Furthermore, Our proposed framework exhibits superior tolerance across the entire range of pose perturbations compared to existing methods. Even in the presence of significant perturbation, the degradation in EF-3DGS’s novel view synthesis fidelity and trajectory estimation performance is substantially lower than that of the other methods.
Figure 8: Qualitative Results of the effectiveness of $\mathcal{L}_{EGM}$.
TABLE V: Effect of each component in EF-3DGS. Evaluation is conducted on 1FPS scenes of Tanks and Temples. The best results are highlighted in bold.
$\mathcal{L}_{EGM}$ $\mathcal{L}_{LEGM}$ $\mathcal{L}_{PBA}$ Fixed GS NVS Pose
[…]
20.53 0.65 0.46 0.1057 0.9768 0.8972
22.16 0.68 0.42 0.0651 0.7529 0.5779
21.07 0.67 0.44 0.0830 0.8869 0.7231
20.96 0.65 0.46 0.0938 0.9875 0.9112
22.83 0.68 0.40 0.0523 0.6387 0.3981
23.46 0.70 0.37 0.0523 0.6387 0.3981
23.09 0.70 0.38 0.0487 0.6259 0.3753
23.96 0.72 0.36 0.0487 0.6259 0.3753
Figure 9: Qualitative Results of the effectiveness of $\mathcal{L}_{LEGM}$.
Figure 10: Qualitative Results of the effectiveness of $\mathcal{L}_{PBA}$.
Figure 11: Robustness of different methods to pose disturbance.
TABLE VI: Effect of the number of event subintervals N.
[…]
$N=2$ 23.56 30+ 1.3h
$N=3$ 24.43 30+ 3h
$N=6$ 24.81 30+ 7h
[…]
CF-3DGS 22.08 30+ 3h
TABLE VII: Comparison of training time vs PSNR for various methods
V-E Ablation Studies
To further investigate the effectiveness of each key component in our proposed EF-3DGS framework, we conduct ablation experiments under the 1 FPS setting on the Tanks and Temples dataset. The results are presented in Table V.
V-E1 Effect of $\mathcal{L}_{EGM}$
Analysis of Table V demonstrates that removing all components substantially degrades both rendering quality and pose estimation accuracy. This finding underscores the significant challenges faced by the original CF-3DGS in high-speed scenarios. Fig. 8 further illustrates this issue, revealing that while CF-3DGS can produce high-quality renderings at training views, it struggles with adjacent novel views. Moreover, the rendered depth maps expose considerable inaccuracies in CF-3DGS’s scene geometry reconstruction. The incorporation of $\mathbf{\mathcal{L}_{EGM}}$ yields marked improvements in both rendering quality and pose estimation precision, as evidenced by Table V and Fig. 8. This experimental outcome suggests that $\mathbf{\mathcal{L}_{EGM}}$ effectively harnesses the continuous brightness change data captured by event cameras, thereby compensating for the inter-frame loss of scene brightness information. Consequently, the 3DGS training optimization achieves multi-view geometric consistency. Notably, among all components, $\mathbf{\mathcal{L}_{EGM}}$ demonstrates the most profound impact on overall method performance, establishing its role as a fundamental component of our approach.
V-E2 Effect of $\mathcal{L}_{LEGM}$
As demonstrated in Table V, $\mathbf{\mathcal{L}_{LEGM}}$ further enhances both rendering quality and pose estimation accuracy when applied in conjunction with $\mathbf{\mathcal{L}_{EGM}}$. Fig. 9 provides a visual illustration of $\mathbf{\mathcal{L}_{LEGM}}$’s efficacy. Ideally, with accurate depth and pose estimation, the Image of Warped Events (IWE) should resemble the reference image. However, in the absence of $\mathbf{\mathcal{L}_{LEGM}}$, the generated IWE exhibits blurring effects, indicative of inaccurate pose estimation. More critically, as highlighted in the boxed area of Fig. 9, the presence of self-similar regions with repetitive textures poses additional challenges for scene reconstruction and pose estimation. Significant errors in scene reconstruction leading to inaccurate depth estimates further compromise the inherent structure of the IWE. The introduction of $\mathbf{\mathcal{L}_{LEGM}}$ allows for the simultaneous reflection of both pose estimation and scene reconstruction errors in the IWE. By optimizing the sharpness of the IWE, $\mathbf{\mathcal{L}_{LEGM}}$ effectively improves both rendering quality and pose estimation precision.
V-E3 Effect of $\mathcal{L}_{PBA}$ and Fixed-GS Strategy
Comparing the last three rows of Table V, we observe that both $\mathbf{\mathcal{L}_{PBA}}$ and Fixed-GS strategies significantly enhance rendering quality. $\mathbf{\mathcal{L}_{PBA}}$, by enforcing multi-view consistency, slightly improves pose estimation accuracy, while the Fixed-GS strategy, applied solely during scene optimization, has no impact on pose estimation. Fig. 10 visualizes the effects of these components. When both are removed, the reconstructed scene loses almost all color information. This is attributed to the substantially higher number of event frames than image frames, causing event-guided optimization to dominate scene reconstruction and suppress color information from image frames. $\mathbf{\mathcal{L}_{PBA}}$ alone, which reprojects neighboring video frames onto event frames, only partially mitigates this imbalance. The Fixed-GS strategy employs a two-stage scene optimization approach for the separation of the color and structure of the scene. However, this approach does not fully leverage the supervisory signals from adjacent frames. The combination of $\mathbf{\mathcal{L}_{PBA}}$ and Fixed-GS strategies complements each other, more effectively addressing the color distortion issue.
V-E4 Computational Efficiency and Impact of Subinterval Number N
To investigate the impact of the number of subintervals N on our method, we conduct ablation studies on the Tanks and Temples dataset under the 2 FPS setting. The results are shown in Table VII and figure on the right. We test the speed on RTX2080ti. Note that in the t-PSNR figure, circular nodes represent pose-free methods, while diamond-shaped nodes indicate methods that require precomputed poses. The results in Table VII reveal that increasing N improves PSNR, indicating that finer temporal discretization enhances reconstruction quality. However, this improvement comes at the cost of extended training time, increasing from 1.3 hours at N=2 to 7 hours at N=6. Notably, our method maintains real-time rendering performance, thanks to the efficient rendering capabilities of 3DGS. Even at N=2, our approach outperforms baselines in both PSNR and rendering speed. The setting of N=3 offers a good balance, achieving higher PSNR than CF-3DGS with comparable training time. However, it’s important to note that the optimal N may vary for different frame rates and motion characteristics.
Figure 12: Study of the different evalution metrics with respect to $\lambda_{cm}$
Figure 13: Study of the different evalution metrics with respect to $\lambda_{grad}$
Figure 14: Study of the different evalution metrics with respect to $\lambda_{PBA}$
V-F Analysis of Loss Coefficients
Different components of our method play crucial roles in overall performance. To comprehensively understand these effects and determine optimal weight settings, we conducted a detailed analysis of the loss weights: $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$.
V-F1 Contrast Maximization Coefficient $\lambda_{cm}$
As shown in Fig. 14, the contrast maximization coefficient $\lambda_{cm}$ significantly affects both NVS quality and pose estimation. We observe that a value of 0.1 achieves the best overall performance. Lower values (0.01) lead to decreased performance, likely due to insufficient utilization of motion information from events. Higher values (0.5) also result in performance degradation, possibly due to over-reliance on the contrast maximization term at the expense of other constraints. The optimal $\lambda_{cm}$ suggests that while event data is crucial, it should not dominate the reconstruction process. This balance allows our method to leverage the high temporal resolution of events without sacrificing the global supervision provided by frame data.
V-F2 Gradient-based Loss Coefficient $\lambda_{grad}$
As shown in Fig. 14, the gradient-based loss coefficient $\lambda_{grad}$ shows optimal performance at 0.2. Lower values (0.1) slightly decrease performance, while higher values (0.3, 0.5) lead to more significant drops in both NVS quality and pose accuracy. This is likely because $\mathcal{L}_{EGM}$ and 3DGS primarily focus on rendering absolute brightness, and excessive gradient constraints may interfere with this process. Therefore, a moderate $\lambda_{grad}$ value is crucial to balance gradient information with the primary rendering objectives.
V-F3 Photometric Bundle Adjustment Coefficient $\lambda_{PBA}$
As illustrated in Figure 14, our method demonstrates remarkable stability across $\lambda_{PBA}$ values ranging from 0.1 to 0.5. However, Fig. 10 illustrates that while $\mathcal{L}_{PBA}$ has minimal effect on PSNR and ATE metrics, its absence leads to noticeable color distortion artifacts in visual results. This underscores the importance of $\lambda_{PBA}$ in maintaining visual fidelity. Conversely, setting $\lambda_{PBA}$ to 1.0 results in significant performance degradation, indicating that overemphasis on cross-view consistency can be counterproductive.
VI Limitations
Our method combines event streams with conventional video frames, which require the input data to be time-ordered. So it can not handle the unordered image input. In high-speed situations, video frames may suffer from severe motion blur, which can degrade the quality of the visual information. Future research could focus on this challenge. And, our method relies on several key parameters (e.g., $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$) that may require manual tuning for optimal performance across different speed scenarios. This dependency on scene-specific parameter settings could limit the method’s adaptability to diverse environments. Future work should explore self-adaptive parameter adjustment strategies to enhance the method’s versatility and ease of use across various reconstruction tasks.
VII Conclusions
In this work, we propose Event-Aided Free-Trajectory 3DGS (EF-3DGS), a novel framework that seamlessly integrates event camera data into the task of reconstructing 3DGS from casually captured free-trajectory videos. Our method effectively leverages the high temporal resolution and motion information encoded in event streams to enhance the 3DGS optimization process, leading to improved rendering quality and accurate camera pose estimation.

### v3

@@
 Abstract
-Scene reconstruction from casually captured videos has wide applications in real-world scenarios.
-With recent advancements in differentiable rendering techniques, several methods have attempted to simultaneously optimize scene representations (NeRF or 3DGS) and camera poses.
-Despite recent progress, existing methods relying on traditional camera input tend to fail in high-speed (or equivalently low-frame-rate) scenarios due to insufficient observations and large pixel displacements between adjacent frames.
+Scene reconstruction from casually captured videos has wide real-world applications.
+Despite recent progress, existing methods relying on traditional cameras tend to fail in high-speed scenarios due to insufficient observations and inaccurate pose estimation.
 Event cameras, inspired by biological vision, record pixel-wise intensity changes asynchronously with high temporal resolution and low latency, providing valuable scene and motion information in blind inter-frame intervals.
-In this paper, we introduce the event camera to aid scene construction from a casually captured video for the first time, and propose Event-Aided Free-Trajectory 3DGS, called EF-3DGS, which seamlessly integrates the advantages of event cameras into 3DGS through three key components.
-First, we leverage the Event Generation Model (EGM) to fuse events and frames, supervising the rendered views observed by the event stream.
-Second, we adopt the Contrast Maximization (CMax) framework in a piece-wise manner to extract motion information by maximizing the contrast of the Image of Warped Events (IWE), thereby calibrating the estimated poses.
-Besides, based on the Linear Event Generation Model (LEGM), the brightness information encoded in the IWE is also utilized to constrain the 3DGS in the gradient domain.
-Third, to mitigate the absence of color information of events, we introduce photometric bundle adjustment (PBA) to ensure view consistency across events and frames.
-Additionally, we propose a Fixed-GS training strategy that separates the optimization of scene structure and color, effectively addressing color distortions caused by the lack of color information in events.
+In this paper, we introduce the event cameras to aid scene construction from a casually captured video for the first time, and propose Event-Aided Free-Trajectory 3DGS, called EF-3DGS, which seamlessly integrates the advantages of event cameras into 3DGS through three key components.
+First, we leverage the Event Generation Model (EGM) to fuse events and frames, enabling continuous supervision between discrete frames.
+Second, we extract motion information through Contrast Maximization (CMax) of warped events, which calibrates camera poses and provides gradient-domain constraints for 3DGS.
+Third, to address the absence of color information in events, we combine photometric bundle adjustment (PBA) with a Fixed-GS training strategy that separates structure and color optimization, effectively ensuring color consistency across different views.
 We evaluate our method on the public Tanks and Temples benchmark and a newly collected real-world dataset, RealEv-DAVIS.
-Our method achieves up to 2dB higher PSNR and 40% lower Absolute Trajectory Error (ATE) compared to state-of-the-art methods under challenging high-speed scenarios.
-Our project page is https://lbh666.github.io/ef-3dgs/.
-Index Terms:
-Event Camera, Novel View Synthesis, 3D Gaussian Splatting, Neural Rendering
+Our method achieves up to 3dB higher PSNR and 40% lower Absolute Trajectory Error (ATE) compared to state-of-the-art methods under challenging high-speed scenarios.
 Figure 1: Free-trajectory 3DGS under high speed.
@@
 The colored dots (red for positive and blue for negative events) in the top row represent the event data.
-Given free-trajectory video and corresponding event stream under the high-speed scenario, we aim to jointly estimate poses and the 3DGS.
+We leverage continuous event streams to aid discrete video frames captured along free trajectories in high-speed scenarios, jointly optimizing camera poses and reconstructing the 3DGS.
 Our method surpasses current state-of-the-art methods in terms of both rendered results (middle row) and pose estimation (bottom row).
-I Introduction
+1 Introduction
 In recent years, Neural Radiance Fields (NeRF) [30, 1, 31] and 3D Gaussian splatting (3DGS) [20, 9] have made significant progress in novel view synthesis tasks.
 Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering.
-While subsequent methods [2, 3, 6, 7] have made great strides in rendering quality and speed, how to reconstruct the scene from a video captured with free camera trajectories remains an open problem.
-It has great application potential in the real world, such as in VR/AR, video stabilization, and mapping.
+While subsequent methods [1, 31, 7, 2] excel with posed images, reconstructing scenes from videos with free camera trajectories remains challenging despite its applications in VR/AR, video stabilization, and mapping.
 To tackle this challenging task, several efforts have been made.
-Free-trajectory videos often cover a vast field of view, so a key challenge is how to preserve the detailed textures of the foreground while representing the overall large-scale scene.
-Previous methods [8, 9] adopt different spatial transformation and allocation approaches based on neighboring frames’ viewing angles and range to achieve efficient and high-quality scene reconstruction.
-Another challenge is that accurate pose estimation is often difficult to obtain in free-trajectory scenarios.
-Previous methods [9, 10] have incorporated priors, such as the depth [11, 12, 13] and the optical flow [14], into their loss functions to improve pose estimation or modify their optimization strategies.
-While these methods above can render photo-realistic images in typical free-trajectory scenarios, the rendering quality degrades significantly in high-speed scenarios (or equivalent low-frame-rate scenarios).
-This type of scenario has essential applications in the real world, such as autonomous driving or FPV exploration.
-Besides, the significant motion between consecutive frames poses challenges for feature matching and further abrupt pose estimation (as shown in the third row of Fig.1).
-This makes it even more crucial for the algorithm to robustly estimate the poses in such scenarios.
+Accurate pose estimation is often difficult to obtain in free-trajectory scenarios, which directly impacts the quality of scene reconstruction.
+One line of work draws inspiration from Simultaneous Localization and Mapping (SLAM).
+They [28, 9, 19] follow its optimization paradigm, progressively optimizing camera trajectories and alternating between camera pose and scene refinement.
+Another line of work [4, 19, 25, 9, 42] explores incorporating additional geometric or motion priors such as depth estimation [29, 36] or optical flow [41] to establish constraints beyond photometric rendering loss.
+While these methods can render photo-realistic images in typical free-trajectory scenarios, both their rendering quality and pose estimation accuracy degrade significantly in high-speed scenarios (or equivalently low-frame-rate scenarios) as shown in Fig.
+1.
+Such high-speed scenarios have essential applications such as autonomous driving and First-Person View (FPV) exploration.
 The performance degradation of prior methods can be attributed to two primary factors.
-Firstly, the limited number of camera observations leads to an under-constrained scene reconstruction problem.
+First, the limited number of camera observations leads to an under-constrained scene reconstruction problem.
 This can cause the scene representation to converge to a trivial solution [32, 47, 42], where the model overfits to the training views without capturing the correct underlying geometry structure.
-Consequently, the synthesized images from novel viewpoints exhibit visual artifacts.
-Secondly, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which was leveraged by previous techniques such as LocalRF [9] and NoPe-NeRF [10].
-This significant violation greatly exacerbates the ill-posedness of the joint optimization of scene and camera poses, further deteriorating the quality of the synthesized novel views.
+Second, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which is leveraged by previous methods.
+Moreover, geometric and motion priors like optical flow and feature matching become unreliable in such scenarios.
+These significant violations greatly exacerbate the ill-posedness of the joint optimization of scene and camera poses.
 Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur [12, 17, 45, 46, 52, 6].
 The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames.
-Moreover, in scenarios with stable scene illumination, the event data naturally encodes the motion information of the scene [24, 25, 26], containing rich motion cues.
-These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios [27].
+Moreover, the event data naturally encodes the motion information of the scene [24, 40, 44], containing rich motion cues.
+These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios.
 However, seamlessly integrating the aforementioned benefits of event cameras is nontrivial.
-First, NeRF or 3DGS renders absolute pixel brightness, which aligns with image data.
-In contrast, event stream records sparse scene brightness changes, lacking the ability to perceive absolute brightness, and most event cameras lack color information.
-Second, the motion cues in event cameras is encoded in the spatio-temporal correlations of the event stream.
-In contrast, previous methods typically sample and optimize frames and pixels independently, without explicitly considering these temporal correlations.
-In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage the high temporal resolution property of event cameras.
-Specifically, based on the Event Generation Model (GEM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream.
-This allows us to effectively utilize the brightness cues encoded in the event streams between adjacent frames, providing rich supervisory signals to alleviate the problem of insufficient sparse viewpoints.
-Regarding the pose estimation challenge, we introduce the CMax [28] framework to exploit the spatio-temporal correlations of events.
-We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive time.
+First, 3DGS renders absolute pixel brightness, which aligns with image data.
+Event cameras, however, record sparse differential brightness changes.
+Directly integrating the differential operations into 3DGS may amplify noise and lead to ill-conditioned optimization problems with high sensitivity to parameter initialization and perturbations.
+Second, event cameras encode motion through continuous spatio-temporal trajectories of events.
+In contrast, frame-based data inherently discretizes continuous motion, forcing traditional methods to rely on correspondence matching, which fails in high-speed scenarios with large inter-frame displacements.
+These fundamental challenges require carefully designed method that bridges the gap between the event data and 3DGS optimization.
+In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage its high temporal resolution property.
+Our approach comprises three key components: (1) In the Event Generation Model (EGM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream.
+This allows us to utilize the brightness cues encoded in the event stream between adjacent frames, providing rich supervisory signals to alleviate the insufficient sparse view issues.
+(2) In the Linear Event Generation Model (LEGM), regarding the pose estimation challenge, we introduce the CMax [11] framework to exploit the spatio-temporal correlations of events.
+We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive frames.
 We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns.
-Furthermore, through the Linear Event Generation Model (LEGM) [29, 30], we establish a connection between motion and brightness changes.
+Furthermore, through the Linear Event Generation Model [10, 13], we establish a connection between motion and brightness changes.
 This allows us to constrain the 3DGS in the gradient domain using the IWE.
-Finally, as event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this.
+(3) As most event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this.
 PBA recovers color by optimizing reprojection errors onto RGB frames, while Fixed-GS enables separate optimization of scene structure and color.
@@
   • Experiments on both public benchmarks and real-world datasets demonstrate that our method significantly outperforms existing state-of-the-art approaches in terms of both rendering quality and trajectory estimation accuracy.
-Our method achieves up to 2dB higher PSNR and 40% lower Absolute Trajectory Error compared to state-of-the-art methods in high-speed scenarios.
-II RELATED WORKS
-II-A Large Scale Scene Reconstruction
-The reconstruction of large scenes has emerged as a significant area of research over the past decade.
-Traditional approaches [31, 32] rely on Structure from Motion techniques to recover camera poses and sparse 3D point clouds.
-The introduction of NeRF [1] marked a paradigm shift in scene representation and reconstruction.
-Several works [33, 34, 2] have attempted to extend NeRF to unbounded scenes.
-F2-NeRF [8] adopts perspective warping based on the local viewing range to handle free-viewpoint unbounded scenes.
-To extend unbounded scenes to large-scale scenes, Mega-NeRF [35] and Block-NeRF [36] adopt a divide-and-conquer strategy.
-LocalRF [9] allocates new radiance fields dynamically according to the distance between the current view and the previous radiance field.
-Recently, 3D Gaussian Splatting (3DGS) has been extended to large-scale scene reconstruction due to its efficient rendering capabilities and explicit 3D representation.
-CityGaussian [37] and VastGaussian [38] adopt spatial allocation strategies similar to Mega-NeRF [35].
-CityGaussian [37] dynamically selects poses for optimization based on visibility criteria, while VastGaussian [38] employs a visibility-driven progressive approach to dynamically select both training poses and Gaussian points for optimization.
-In contrast to these methods, which primarily focus on scalability and often require precomputed camera poses from COLMAP [39], we focus on the challenges of reconstructing scene from casually captured free-trajectory videos without relying on precomputed poses.
-II-B Joint Pose and Scene Optimization
+2 Related Works
+2.1 Joint Pose and Scene Optimization
 The research community has recently focused on developing methods [42, 28, 50, 8, 4, 23, 9] that can be optimized without requiring precomputed camera poses.
@@
 Another line of work has investigated incorporating additional constraints to make the problem more tractable.
-SPARF [17] incorporates cross-view feature matching constraints and depth consistency regularization.
-More recent approaches [10, 9, 5] leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation, to obtain additional priors and better address the pose-free NeRF reconstruction problem.
-Exploiting 3DGS’s explicit representation, CF-3DGS [5] directly back-projects Gaussian points using depth maps, enhancing the utilization of depth priors.
-While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed ( or equivalently sparse-view) scenarios or rely on a good pose initialization.
+LocalRF [28] leverages the prior assumption of continuous motion between adjacent frames and progressively adds and optimizes camera poses.
+More recent approaches [4, 28, 9] leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation.
+Exploiting 3DGS’s explicit representation, CF-3DGS [9] directly back-projects Gaussian points using depth maps.
+While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed scenarios or rely on a good pose initialization.
 Our approach addresses these issues by leveraging motion and brightness cues from event streams.
-II-C Event-Based Novel view Synthesis
-Recent works have explored the integration of event cameras into the NeRF framework.
-Early approaches, such as E-NeRF [43] and EventNeRF [44], utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes calculated by the accumulated events.
+2.2 Event-Based Novel View Synthesis
+Recent works have explored the integration of event cameras into the NeRF or 3DGS framework.
+Early approaches, such as E-NeRF [21] and EventNeRF [37], utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes.
 Building upon this, Robust e-NeRF [26] incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise.
 Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods.
-For instance, E2NeRF [46] and EvDeblurNeRF [47] leverage the Event Double Integral (EDI) [48] model to address the deblurring problem, while DE-NeRF [49] and EvDNeRF [50] leverage the high temporal resolution property of event camera to capture fast-moving elements in dynamic scene.
-A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we jointly optimize for both the camera poses and the scene representation.
-Furthermore, while previous works have been limited to simulated environments or simple indoor scenes, we evaluate our approach in large-scale outdoor scenarios.
-III Preliminary
+For instance, E2NeRF [35] and EvDeblurNeRF [5] leverage the Event Double Integral (EDI) [33] model to address the deblurring problem, while DE-NeRF [27] and EvDNeRF [3] leverage the high temporal resolution property of event cameras to capture fast-moving elements in dynamic scene.
+More recently, Event-3DGS [16] and EaDeblur-GS [51] have extended previous approaches to 3D Gaussian Splatting, achieving superior rendering quality and real-time performance.
+A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we target free-trajectory scenarios, jointly optimizing for both the camera poses and the scene representation.
+Furthermore, while previous works have been limited to simulated and simple environments, we evaluate our approach in large-scale outdoor scenarios with complex motions and lighting conditions.
+Figure 2: Method overview.
+Our method takes video frames and event stream as input.
+During the first stage, we progressively add new event images, leveraging the events and most recent frame to establish the event-driven optimization.
+During the second stage, we adopt the Fixed-GS strategy to mitigate the color distortion of 3DGS.
+The details of $\mathcal{L}_{LEGM}$ and CMax framework are illustrated in Fig.
+3.
+3 Preliminary
 3DGS [20] parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information.
 Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$.
-The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$,
-$G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ (1)
-To preserve the valid positive semi-definite property during optimization, the $\mathbf{\Sigma}$ is derived by the scale factor $s$ and the rotation factor $R$,
-$\mathbf{\Sigma}=RSS^{T}R^{T}.$ (2)
-Before rendering, the Gaussian is first projected to the 2D image plane:
-$\mathbf{\Sigma}^{2D}=JW\mathbf{\Sigma}W^{T}J^{T},$ (3)
-where J is the Jacobian of the affine approximation of the projective transformation, and W is the viewing transformation.
-During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending.
+The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$, $G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending.
 The color of each pixel is calculated via blending N ordered overlapping points:
@@
 where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel.
-With the forward rendering procedure, we can optimize 3DGS by minimizing the photometric loss between observation and rendered pixels:
-$\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ (5)
-where the first term is a standard $\mathcal{L}_{1}$ loss between the target and rendered images, $\mathcal{L}_{D-SSIM}$ is a differentiable version of structural similarity index and $\lambda$ is balancing weight which is set to 0.2 following [4].
-As this optimization is highly sensitive to local minima, 3DGS additionally employs periodic adaptive densification and pruning of splats.
-By integrating depth $t_{i}$ of each Gaussian in Equation 4 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
-$\hat{D}(\mathbf{r})=\sum_{i=1}^{N}t_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}).$ (6)
-Figure 2: Method Overview.
-Our method takes video frames $\{I_{i}\}$ and event stream $\varepsilon=\{\mathbf{e}_{k}\}$ as input.
-During training, we random sample $t\in\{t_{i,j}\}$, leveraging the events within current subinterval and most recent frame to establish the proposed three main constrains,$\mathcal{L}_{EGM}$, $\mathcal{L}_{LEGM}$ and $\mathcal{L}_{PBA}$.
-The colored dots (red for positive and blue for negative events) represent the event data.
-The details of $\mathcal{L}_{LEGM}$ and CMax framework are illustrated in Fig.
-3
-IV Method
-In this section, we introduce our method.
+With the forward rendering procedure, we can optimize 3DGS by minimizing a weighted combination loss of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ between observation and rendered pixels: $\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ where $\lambda$ is balancing weight which is set to 0.2 following [20].
+By integrating depth $d_{i}$ in Equation 1 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
+$\hat{D}(\mathbf{r})=\sum_{i=1}^{N}d_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}).$ (2)
+4 Method
 The overall framework is shown in Fig.
 2.
-Given a video of a free-trajectory $\{I_{i}\}$ captured at time $\{t_{i}\}$ and an event stream $\varepsilon=\{\mathbf{e}_{k}\}$, our goal is to reconstruct the 3DGS of the scene and the corresponding camera trajectory$\{T_{i}\}$.
-The photometric loss of the original 3DGS follows an analysis-by-synthesis process.
-Similarly, in order to incorporate the events into the 3DGS optimization process, we delve into the imaging principles of event cameras.
-Revolving around the Event Generative Model (EGM) and Linearized Event Generative Model (LEGM), we establish constraints from the connection between motion and brightness changes.
-Further, through photometric bundle adjustment (PBA) and Fix-GS training strategy, we mitigate the absence of color information of events and further strengthen the consistency of cross-view rendering.
-IV-A EGM Driven Optimization
-The Event Generation Model (EGM) describes the base working principle of event cameras, which capture pixel-wise brightness changes asynchronously.
+Given a video of a free-trajectory $\{I_{i}\}$ captured at time $\{t_{i}\}$ and the event stream $\varepsilon=\{\mathbf{e}_{k}\}$, our goal is to reconstruct the 3DGS of the scene and the corresponding camera trajectory$\{T_{i}\}$.
+Following the analysis-by-synthesis paradigm of 3DGS, we extend this approach by incorporating event camera data through two fundamental imaging principles: the Event Generative Model (EGM) and Linear Event Generative Model (LEGM).
+To address the absence of color information in events and ensure cross-view consistency, we further introduce photometric bundle adjustment (PBA) and a Fixed-GS training strategy.
+4.1 EGM Driven Optimization
+The Event Generation Model (EGM) describes how event cameras asynchronously record pixel-wise brightness changes.
 When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,
@@
 where $L\doteq log(I)$ is the logarithm of intensity, $p_{k}\in\{-1,+1\}$ indicates the polarity of brightness changes, and $t_{k}$ is the triggered timestamp.
-By leveraging the high temporal resolution of event cameras, we can effectively constrain the 3DGS optimization process in the temporal domain.
-To incorporate the event data into the optimization process, we first divide the time interval between two adjacent video frames $I_{i}$ and $I_{i+1}$ into N smaller subintervals $\varepsilon_{i,j}=\{\mathbf{e}_{k}|t_{i,j}\leq t_{k}\leq t_{i,j+1},\Delta t=\frac{t_{i+1}-t_{i}}{N},t_{i,j}=t_{i}+j\cdot\Delta t\}$.
-This allows us to form event frames at a higher temporal resolution than the original video frames:
+As shown in Fig.
+2 (1.1), to leverage the high temporal resolution of events, we first divide the time interval between two adjacent video frames $I_{i}$ and $I_{i+1}$ into N smaller subintervals $\varepsilon_{i,j}=\{\mathbf{e}_{k}|t_{i,j}\leq t_{k}\leq t_{i,j+1},\Delta t=\frac{t_{i+1}-t_{i}}{N},t_{i,j}=t_{i}+j\cdot\Delta t\}$.
+This allows us to form accumulated event frames at a higher temporal resolution:
 $E_{i,j}=\sum_{e_{k}\in\varepsilon_{i,j}}p_{k}.$ (4)
-We then reconstruct the latent intensity image $I_{t}$ at any intermediate time $t\in\{t_{i,j}\}$ between two video frames $I_{i}$ and $I_{i+1}$ using the accumulated events and the most recent frame:
+We then reconstruct the latent intensity image $I_{t}$ at any intermediate time $t\in\{t_{i,j}\}$ by integrating the accumulated events with the most recent frame:
 $I_{t}=I_{i,j}=\begin{cases}I_{i,0}\cdot\exp(\sum_{n=0}^{j-1}E_{i,n}\cdot C)&\text{if }j>0\\ I_{i,0}&\text{if }j=0\end{cases}.$ (5)
-Then, the event-based rendering loss is introduced to enforce the consistency between rendered and latent intensity images:
+This latent intensity image provides a supervisory signal for our event-based rendering loss:
 $\mathcal{L}_{EGM}=(1-\lambda)\mathcal{L}_{1}(\hat{I}_{t},I_{t})+\lambda\mathcal{L}_{D-SSIM}(\hat{I}_{t},I_{t}).$ (6)
-This loss function ensures that the 3DGS optimization process is constrained not only by the sparse video frames but also by the dense event data, leading to improved reconstruction quality.
-It is worth noting that when $j=0$, the latent intensity image $I_{t}$ is equal to the first most recent video frame $I_{i,0}$ and $\mathcal{L}_{EGM}$ reduces to the standard image rendering loss.
-IV-B Unified CMax and LEGM Optimization
-While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by the events, it does not explicitly exploit the motion information encoded in the event stream.
-To further improve the accuracy of camera poses and the reconstructed 3DGS, we introduce the Contrast Maximization (CMax) [28, 51, 52] framework and the Linear Event Generation Model (LEGM) [29, 30, 53] in this section.
-These models complement the EGM-driven optimization by extracting motion cues and gradient information from the event stream, providing additional constraints for the 3DGS optimization process.
-Under the assumption of constant scene illumination, events are triggered solely by the motion of scene edges.
-The edges in the scene form continuous trajectories in the $(x,y,t)$ space.
-Intuitively, by back-projecting the events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE).
-Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the currently estimated motion.
-This insight motivates us to derive a motion field by leveraging the rendered depth from 3DGS using eq.
+By enforcing consistency between rendered and latent intensity images, this loss effectively utilizes the brightness information encoded in event streams between adjacent frames, addressing the challenge of sparse viewpoints in high-speed scenarios.
+4.2 Unified CMax and LEGM Optimization
+While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by events, it does not explicitly exploit the motion information encoded in the event stream.
+To address this, we introduce the Contrast Maximization (CMax) [11, 34, 38, 15] framework and the Linear Event Generation Model (LEGM) [10, 13, 56].
+These models complement the previous EGM-driven optimization.
+Under constant scene illumination, events are triggered by the motion of scene edges, forming continuous trajectories in (x, y, t) space.
+As shown in Fig.
+2 (1.2), by warping(back-projecting) events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE).
+Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the estimated motion.
+This insight motivates us to derive the motion field by leveraging the rendered depth from 3DGS using Eq.
 2 and the relative camera motion between neighboring timestamps.
-By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the relative poses between adjacent frames.
-Instead of warping each event as in the original CMax framework, we adopt a piece-wise warping approach for efficiency.
-Specifically, we first accumulate the events within each sub-interval to obtain event frames $E_{i,j}$ as in eq.
-8 .
-Then, for randomly sampled timestamps $t_{ref}=t_{i,j}$, we warp the event frames in the neighboring $2r$ sub-intervals to $t_{ref}$:
-$E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),$ (11)
-where $m\in[-r,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow which can be derived from the rendered depth $\hat{D}$ in  eq. 6 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
-$F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),$ (12)
+By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the camera poses.
+As shown in Fig.
+3, for efficiency, we adopt a piece-wise warping approach instead of warping individual events.
+Specifically, for current timestamp $t_{ref}=t_{i,j}$, we warp the event frames from previous $r$ sub-intervals:
+$\vspace{1pt}E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),\vspace{1pt}$ (7)
+where $m\in[0,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow derived from the rendered depth $\hat{D}$ in  Eq. 2 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
+$\vspace{1pt}F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),\vspace{1pt}$ (8)
 where $T_{i,j\rightarrow j-m}=T_{i,j-m}T_{i,j}^{-1}$, $\Pi$ projects a 3D point to image coordinates and $\Pi^{-1}$ unprojects a pixel coordinate and depth into a 3D point.
 Then the image of piece-wise warped events (IPWE) at timestamp $t_{i,j}$ is computed by averaging the warped event frames:
-$\textrm{IPWE}_{i,j}=\frac{1}{2r+1}\sum_{m=j-r}^{j+r}E_{i,m\rightarrow j}\approx\frac{1}{C}\Delta L.$ (13)
+$\textrm{IPWE}_{i,j}=\frac{1}{r+1}\sum_{m=j-r}^{j}E_{i,m\rightarrow j}\approx\frac{1}{C}\Delta L.$ (9)
 Following the Cmax framework, we maximize the variance of the IPWE, which is equivalent to minimize its opposite:
 $\mathcal{L}_{cm}=-\textrm{Var}(\textrm{IPWE}_{i,j}).$ (10)
-Maximizing the contrast encourages the motion field to align with the actual motion of edges, leading to sharper IPWEs.
-This objective helps to improve the accuracy of the estimated camera poses and the geometric consistency of the 3DGS reconstruction.
-Furthermore, based on the Linear Event Generation Model (LEGM) [29, 30], the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of a image):
+Furthermore, based on the Linear Event Generation Model (LEGM) [10, 13], the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of an image):
 $\Delta L(\mathbf{u})=-\nabla L\cdot\dot{\mathbf{u}}\approx L(\mathbf{u})-L(\mathbf{u}+\dot{\mathbf{u}}).$ (11)
@@
 $C\cdot\textrm{IPWE}_{i,j}=\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}).$ (12)
-Based on this relationship, we formulate an additional gradient-based loss to constrain the 3DGS optimization:
-$\mathcal{L}_{grad}=||C\cdot\textrm{IPWE}_{i,j}-(\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}))||^{2},$ (17)
+Note that to compute $F_{i,j\rightarrow j+1}$, we estimate $T_{i,j,j+1}$ by leveraging the assumption of locally linear motion from $T_{i,j,j-1}$ and $T_{i,j,j}$.
+Based on this relationship, we formulate an additional gradient-based loss:
+$\vspace{1pt}\mathcal{L}_{grad}=||C\cdot\textrm{IPWE}_{i,j}-(\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}))||^{2},\vspace{1pt}$ (13)
 where $\hat{L}$ is the logarithm of synthesised image $\hat{I}_{t}$. Finally, the full LEGM loss is defined as:
 $\mathcal{L}_{LEGM}=\lambda_{cm}\mathcal{L}_{cm}+\lambda_{grad}\mathcal{L}_{grad},$ (14)
-where $\lambda_{cm}$ and $\lambda_{grad}$ are the balancing weight coefficients for $\mathcal{L}_{cm}$ and $\mathcal{L}_{grad}$ respectively.
-IV-C Photometric Bundle Adjustment
-The aforementioned event-based constrains, $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$, leverage the brightness change and motion information encoded in the event data to constrain 3DGS.
+where $\lambda_{cm}$ and $\lambda_{grad}$ are the balancing weight.
+4.3 Photometric Bundle Adjustment
+The aforementioned event-based constraints, $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$, leverage the brightness change and motion information encoded in the event data to constrain 3DGS.
 However, as event cameras only record brightness changes and lack color perception, directly applying them to 3DGS optimization may lead to inconsistent color rendering.
 To ensure cross-view consistency of the 3DGS rendering, we introduce the Photometric Bundle Adjustment (PBA) term.
-Specifically, for a randomly sampled timestamp $t\in\{t_{i,j}\}$, we establish the following photometric reprojection error:
+Specifically, as shown in Fig. 2 (1.3), for a randomly sampled timestamp $t\in\{t_{i,j}\}$, we establish the following photometric reprojection error:
 $\mathcal{L}_{PBA}=\sum_{\mathbf{u}\in\mathcal{P}}\sum_{I_{s}\in\mathcal{F}}||I_{s}(\mathbf{u}^{\prime})-\hat{I}(\mathbf{u})||^{2},$ (15)
 where $\mathbf{u}^{\prime}=\Pi(T_{i,j-r\rightarrow j}\Pi^{-1}(x,y,\hat{D}(u)))$ represent the coordinate on target view projected from the pixel $\mathbf{u}$ of source view $I_{s}$, $\mathcal{P}$ denotes the pixel samples of current frame, and $\mathcal{F}$ is the candidates of target video frames.
-We select $\mathcal{F}$ to be the adjacent video frames in consideration of computation costs.
-By minimizing $\mathcal{L}_{PBA}$ over sampled pixels in the event frames and neighboring video frames, we encourage the 3DGS model to produce geometrically and photometrically consistent renderings across events and video frames.
+We select $\mathcal{F}$ to be the nearest previous video frame in consideration of computation costs.
+By minimizing $\mathcal{L}_{PBA}$ across sampled views, we encourage the 3DGS model to produce geometrically and photometrically consistent renderings across events and video frames, thus effectively resolving color inconsistencies inherent in event data.
 Figure 3: The illustration of unified CMax and LEGM optimization.
-We warp nearby event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of piece-wise warped events (IPWE).
-The byproduct IWPE is utilized to establish additional constrain on 3DGS in gradient domain.
-IV-D Fixed-GS Training Strategy
-As demonstrated in Fig.
-10, the $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues.
-The absence of color information in event data poses a significant challenge for scene reconstruction in previous NeRF-based methods.
-NeRF models employ implicit representations [54], utilizing MLPs to store scene color and density information in a coupled manner.
-Consequently, color distortions can affect scene density estimations and vice versa.
-In contrast, 3DGS explicitly stores various attributes independently within Gaussian points.
-This characteristic allows us to address color distortion issues specifically while maintaining other attributes.
-To leverage this advantage, we propose a two-stage Fixed-GS scene optimization strategy. In the first stage, optimization is performed across all timestamps:
-$G_{\theta}^{*}=\underset{\alpha,r,s,f}{\operatorname{argmin}}~{}\mathcal{L}_{event},~{}t\in\{t_{i,j}\},$ (20)
-where $\alpha,r,s,f$ is the opacity, rotation, scale factor and spherical harmonics of Gaussians, and t is the sampled timestamp during training.
-This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions.
-The second stage focuses on recovering accurate color information.
-During this phase, optimization is conducted exclusively on video frames.
-Capitalizing on the explicit representation of Gaussians, we optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed.
-[…]
-The ratio between the first and second stages is empirically set to 4:1.
-This approach allows us to effectively address the color distortion problem while preserving the structural and brightness information obtained from the event data.
-IV-E Overall Training Pipeline
-Assembling all loss terms, we get the overall loss function for events:
-$\mathcal{L}_{event}=\mathcal{L}_{EGM}+\mathcal{L}_{LEGM}+\lambda_{PBA}\mathcal{L}_{PBA},$ (22)
-where $\lambda_{PBA}$ are the weighting factor.
-It is noteworthy that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space.
-We borrow the idea of dynamic scene allocation strategies from LocalRF [9].
-Our overall training pipeline builds upon the pipeline of CF-3DGS [5] while introducing novel components to integrate event camera data for free-trajectory scene reconstruction.
-Further implementation details are provided in the Supplementary Material.
-V Experiments
-TABLE I: Quantitative Evaluations on Tanks and Temples.
-Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol.
-The best results are highlighted in bold.
+We warp previous event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of piece-wise warped events (IPWE).
+The byproduct IPWE is utilized to establish additional constraints on 3DGS.
 Methods Pose-Free Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
@@
 ENeRF $\times$ E+F 23.62 0.73 0.37 22.84 0.70 0.38 21.85 0.69 0.41 20.52 0.66 0.46 18.09 0.60 0.52
+Event-3DGS(E+F) $\times$ E+F 26.32 0.78 0.33 25.37 0.76 0.34 24.59 0.75 0.37 23.44 0.72 0.38 22.41 0.69 0.39
 […]
 […]
-TABLE II: Pose accuracy on Tanks and Temples.
-Note that we use COLMAP poses in Tanks and Temples as the “ground truth”.
-The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100.
-Those methods which require precomputed poses are excluded.
+Table 1: Quantitative evaluations on Tanks and Temples.
+Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol.
+The best results are highlighted in bold.
 Methods Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
@@
 CF-3DGS F 0.0594 0.6981 0.4212 0.0637 0.7128 0.4628 0.0712 0.7531 0.5189 0.0859 0.8074 0.6918 0.1057 0.9768 0.8972
-Ev-Baseline E+F 0.0461 0.5972 0.3419 0.0490 0.6269 0.3766 0.0538 0.6728 0.4261 0.0591 0.7094 0.4860 0.0657 0.7597 0.5534
+EvCF-3DGS E+F 0.0461 0.5972 0.3419 0.0490 0.6269 0.3766 0.0538 0.6728 0.4261 0.0591 0.7094 0.4860 0.0657 0.7597 0.5534
 EF-3DGS(Ours) E+F 0.0391 0.5427 0.2885 0.0407 0.5521 0.3064 0.0426 0.5796 0.3271 0.0449 0.5953 0.3671 0.0487 0.6259 0.3753
-Figure 4: Qualitative comparison for novel view synthesis on Tanks and Temples. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
-Figure 5: Pose Estimation Comparison on Tanks and Temples.
+Table 2: Pose accuracy on Tanks and Temples.
+We use COLMAP poses in Tanks and Temples as the “ground truth”.
+The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100.
+Those methods that require precomputed poses are excluded.
+4.4 Fixed-GS Training Strategy
+The $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues.
+To further address this challenge, we propose a two-stage Fixed-GS scene optimization strategy that takes advantage of 3DGS’s explicit attribute representation.
+In the first stage, all the parameters are optimizable and the optimization is performed across all timestamps:
+$G_{\theta}^{*},T_{i,j}^{*}=\underset{\mu,\alpha,r,s,f,T_{i,j}}{\operatorname{argmin}}~{}\mathcal{L}_{event},~{}t\in\{t_{i,j}\},$ (16)
+where $\mu,\alpha,r,s,f$ is the position, opacity, rotation, scale factor and spherical harmonics of the Gaussians, and t is the sampled timestamp during training.
+This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions due to the dominant colorless event supervision overwhelming the sparse RGB frame color supervision.
+The second stage focuses on recovering accurate color information.
+During this phase, optimization is conducted exclusively on video frames.
+We optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed:
+[…]
+The ratio between the first and second stages is empirically set to 4:1.
+This approach allows us to effectively address the color distortion problem while preserving the structural and brightness information obtained from the event data.
+4.5 Overall Training Pipeline
+Assembling all loss terms, we get the overall loss function:
+$\mathcal{L}_{event}=\mathcal{L}_{EGM}+\mathcal{L}_{LEGM}+\lambda_{PBA}\mathcal{L}_{PBA},$ (18)
+where $\lambda_{PBA}$ are the weighting factor.
+Note that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space.
+We incorporate dynamic scene allocation strategies from LocalRF [28] for handling extended video sequences.
+Our overall training pipeline builds upon the progressive optimization scheme of CF-3DGS [9] while introducing novel components to integrate event stream data for robust free-trajectory scene reconstruction.
+Please refer to the Supplementary Material for the algorithm pipeline and additional implementation details.
+5 Experiments
+Figure 4: Qualitative comparison for novel view synthesis.
+The first two rows are from Tanks and Temples and the last row is from RealEv-DAVIS.
+Our approach produces more realistic rendering results with fine-grained details.
+Better viewed when zoomed in.
+Figure 5: Pose estimation comparison.
 We visualise the trajectory (3D plot) and $\textrm{RPE}_{r}$ (color bar) of each method.
 We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
-V-A Dataset
-V-A1 Tanks and Temples
+5.1 Dataset
+Tanks and Temples.
 We conduct comprehensive experiments on the Tanks and Temples dataset [22].
-Similar to LocalRF [9], we adopt the same 9 scenes, covering large-scale indoor and outdoor scenes.
+Similar to LocalRF [28], we adopt 9 scenes, covering large-scale indoor and outdoor scenes.
 For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance.
-Following LocalRF [9], we spatially downsample the videos to a resolution of 1080p.
-Increasing motion speed makes frames further apart, hence it is effectively equivalent to reducing the frame rate.
-We employ various temporal downsamplings to evaluate the robustness of different methods to varying camera speed scenes.
-In the 6 FPS setting, we retain one in every five frames as in LocalRF [9].
-We perform corresponding frame downsampling for the other frame rate settings (4 FPS, 3 FPS, 2 FPS, and 1 FPS).
+Following LocalRF [28], we apply 4× spatial downsampling to the videos.
+To evaluate the robustness under varying camera speeds, we employ varying temporal downsampling of 6 FPS, 4 FPS, 3 FPS, 2 FPS, and 1 FPS.
+The reduction in frame rate effectively creates larger inter-frame displacements, simulating high-speed scenarios.
 To synthesize realistic event data, we first upsample the original videos by [18] and then apply the simulator V2E [14].
-V-A2 RealEv-DAVIS
-Due to the current lack of datasets featuring free-trajectory captures from event cameras, we introduce the RealEv-DAVIS dataset, encompassing four outdoor scenes.
-We use the DAVIS346 camera, which simultaneously records image frames and events at a resolution of 346x260 pixels.
-To simulate real-world free-trajectory scenarios, we capture handheld footage at 25 FPS, with each sequence lasting 40 seconds.
-For SLOW scenarios, we downsample the video clips by keeping every second frame.
-We further reduce the frame rate in FAST scenarios by retaining only one frame out of every five.
-V-B Implementation details
-Our implementation is primarily based on the PyTorch framework, and we follow the optimization parameters by the configuration outlined in the 3DGS [4] unless otherwise specified.
-All the parameters are optimized using Adam [58] with $\beta_{1}=0.9$ and $\beta_{2}=0.99$.
+RealEv-DAVIS.
+Due to the lack of free-trajectory event camera datasets, we introduce RealEv-DAVIS, comprising four outdoor scenes.
+Using a DAVIS346 camera that simultaneously captures frames and events at 346×260 resolution, we record 40-second handheld sequences at 25 FPS.
+We employ COLMAP for ground-truth poses.
+For SLOW scenarios, we retain every second frame, while for FAST scenarios, we keep only one frame per five frames.
+Further details are provided in the supplementary material.
+5.2 Implementation details
+We follow the optimization parameters by the configuration outlined in the 3DGS [20].
 We optimize the camera poses in the representation of quaternion rotation.
-The initial learning rate is set to $10^{-5}$ and gradually decay to $10^{-6}$ until convergence.
-The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5, respectively, based on subsequent experiment results.
+The initial learning rate is set to $10^{-5}$ and gradually decays to $10^{-6}$ until convergence.
+The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5.
 For the division of events between adjacent frames, we maintain a constant interval of $\frac{1}{6}$s for Tanks and Temples and $\frac{1}{25}$s for RealEv-DAVIS, setting the number of subinterval N accordingly.
 For example, in Tanks and Temples, N equals 2 for 3FPS and 6 for 1FPS.
-This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for following CMax warping.
+This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for the following CMax warping.
 The intervals of neighboring warping $r$ in CMax are set to 3.
-That is, we warp the three adjacent event frames to the current sampled timestamp.
-V-C Experimental Setup
-V-C1 Metrics
+The contrast threshold $C$ is set to 0.25 for Tanks and Temples and 0.21 for RealEv-DAVIS.
+We provide detailed ablation studies on these hyperparameters and additional implementation details in the supplementary material.
+5.3 Experimental Setup
+Metrics.
 We evaluate all the methods from two aspects: novel view synthesis and pose estimation.
-For the novel view synthesis task, we report the standard metrics PSNR, SSIM [59], and LPIPS [60] to quantitatively assess the quality of the rendered images.
-To measure the accuracy of pose estimation, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [61, 62], as delineated in [10].
-Given that $\text{RPE}_{r}$ and $\text{RPE}_{t}$ measure relative pose errors between adjacent frames, they are inherently influenced by the frame rate.
-To facilitate a fair comparison across different frame rate settings, we upsample the estimated poses to the same frame rate before evaluation.
-This enables a consistent comparison of relative pose estimation errors across varying frame rate configurations.
-TABLE III: Quantitative Evaluations on RealEv-DAVIS.
-We select the top-performing methods from the previous evaluation.
-Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol.
-The best results are highlighted in bold.
-Methods Pose-Free Input SLOW FAST
-hall building corner outdoor hall building corner outdoor
+For the novel view synthesis task, we report the standard metrics PSNR, SSIM [49], and LPIPS [54].
+For the pose estimation task, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [55, 39], as delineated in [4].
+Since these metrics are inherently influenced by frame rate, we upsample all estimated poses to a consistent temporal resolution before evaluation for fair comparison across different frame rate settings.
+Baselines.
+For a fair comparison, we focus on two categories of methods: (1) For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF [28] and F2-NeRF [48].
+We also include pose-free methods like Nope-NeRF [4] and CF-3DGS [9].
+(2) For event-frame hybrid methods, we consider approaches that fuse events and frames, including ENeRF [21], EvDeblurNeRF [5] and Event-3DGS [16].
+Since no existing method integrates events for free-trajectory scenarios, we implement EvCF-3DGS as a competitive baseline that leverages an event-based frame interpolation network (Time Lens [43]) to temporally upsample frames before feeding them into CF-3DGS.
+5.4 Experimental Results
+We select every ten frames as a test image for NVS evaluation following LocalRF [28].
+Since the camera poses are unknown in our setting, we need to estimate the poses of test views.
+As in iNeRF [53], we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
+Results on Tanks and Temples.
+Tables 1 and 2 demonstrate two key findings: (1) Our event-aided approach achieves up to 3dB higher PSNR and nearly 40% lower trajectory error at 1FPS compared to frame-based methods, indicating the critical value of event data in high-speed scenarios.
+(2) Our method maintains 1.55dB PSNR advantage over Event-3DGS at 1FPS, confirming that our integration framework effectively exploits the nature of event data beyond merely using it.
+Fig.
+4 shows our method produces sharper edges and finer textures, while Fig.
+5 illustrates we achieve more accurate trajectory estimation.
+Methods Input SLOW FAST
+NVS Pose NVS Pose
 […]
-[…]
-[…]
-EvDeblurNeRF $\times$ E+F 20.24 0.5762 19.35 0.5649 20.47 0.5757 22.36 0.7087 18.23 0.5043 16.41 0.4624 17.82 0.4959 19.44 0.6448
-ENeRF $\times$ E+F 21.89 0.5933 21.94 0.5918 22.01 0.5918 23.91 0.7266 20.09 0.5261 18.99 0.4903 19.61 0.5173 20.00 0.6494
-[…]
-[…]
-TABLE IV: Pose accuracy on RealEv-DAVIS.
-The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100.
-Those methods which require precomputed poses are excluded.
-Methods Input SLOW FAST
-hall building corner outdoor hall building corner outdoor
-$\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$ $\textrm{RPE}_{t}$$\downarrow$ $\textrm{RPE}_{r}$$\downarrow$
-LocalRF F 4.4017 2.5067 4.6186 1.9409 1.9296 2.5183 3.4307 1.3005 5.6061 3.2265 7.0611 2.8722 3.4767 3.8950 4.7229 1.8595
-CF-3DGS F 2.7106 0.6035 3.0374 3.2761 1.9437 1.4954 2.2747 0.8053 3.5025 0.7639 4.6758 4.5256 3.4950 2.2855 3.0551 1.1198
-Ev-Baseline E+F 0.7063 0.4386 3.4225 1.1717 2.1283 1.3361 0.8647 0.3455 0.8898 0.7734 5.0707 1.6997 3.6785 2.1552 1.1645 0.5001
-EF-3DGS(Ours) E+F 0.5262 0.3251 2.8008 0.4633 1.8386 1.6168 0.4726 0.3699 0.5424 0.3464 3.5848 0.5589 2.5396 2.2157 0.5218 0.4462
-V-C2 Baselines
-For a fair comparison, we focus on two categories of methods: those solely based on frames and those leveraging both events and frames.
-For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF and F2-NeRF, although the latter requires pre-estimated poses.
-We also include pose-free methods like Nope-NeRF and the current state-of-the-art CF-3DGS.
-To address the out-of-memory (OOM) issues CF-3DGS faces with long video sequences, we implement a slight modification, segmenting long videos into subsequences for individual reconstruction.
-For event-based methods, consistent with EvDeblurNeRF [47], we consider approaches that fuse events and frames.
-This category includes ENeRF [43] and the SOTA deblurring method EvDeblurNeRF [47].
-These hybrid approaches leverage the complementary strengths of both data modalities.
-To the best of our knowledge, no existing method has exploit event cameras to address free-trajectory problem.
-To facilitate a fair comparison, we construct an event-based baseline method built upon CF-3DGS [5].
-It leverages the high temporal resolution property of event cameras by utilizing an event-based frame interpolation network Time Lens [63] to interpolate the original video.
-Then, the temporally upsampled video frames are subsequently fed into CF-3DGS.
-V-D Experimental Results
-To evaluate the results on synthesized novel views, we select every ten frames as a test image following LocalRF [9].
-Since the camera poses are unknown in our setting, we need to estimate the poses of test views.
-As in NeRFmm [40] and iNeRF [64], we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
-V-D1 Results on Tanks and Temples
-Quantitatively, the experimental results in Table I and Table II demonstrate the superiority of event-based methods (EF-3DGS and Ev-Baseline) over frame-based baselines in both pose estimation accuracy and rendering quality.
-This highlights the potential of event cameras for free-trajectory scene reconstruction.
-Furthermore, our method surpasses Ev-Baseline across all evaluation metrics under varying frame rate settings, indicating that our proposed method can better effectively integrate property of the high temporal resolution and motion information encoded in the event streams into the scene reconstruction process.
-Qualitatively, as shown in Fig.
-4, our method generates more realistic and detailed renderings compared to the baseline.
-The synthesized views of our method exhibit sharper edges and better preservation of finer textures.
-In contrast, the baseline methods suffer from various artifacts, such as blurriness and texture inconsistencies.
-Moreover, the visualization of estimated camera trajectories in Fig.
-5 indicates that our estimated trajectory is better aligned with the ground truth than other methods, and the $\textrm{RPE}_{r}$ is significantly smaller.
-Figure 6: Qualitative comparison for novel view synthesis on RealEv-DAVIS. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
-Figure 7: Pose Estimation Comparison on RealEv-DAVIS.
-We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (color bar) of each method.
-We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
-V-D2 Results on RealEv-DAVIS
-In Table III and Table IV, we report results on the self-collected dataset, RealEv-DAVIS.
-We select the top-performing methods from the previous evaluation.
-EF-3DGS demonstrates robust generalization to real-world scenarios, despite challenges such as sensor noise absent in simulated environments.
-Notably, the performance gap widens in FAST scenarios, with PSNR improvements of nearly 2 dB over baselines, emphasizing the method’s efficacy in handling high-speed motion in real-world settings.
-Qualitative results in Fig.
+LocalRF F 20.83 0.6074 3.60 2.07 17.62 0.5192 5.22 2.96
+CF-3DGS F 22.68 0.6287 2.49 1.55 17.59 0.5204 3.68 2.17
+EvDeblurNeRF E+F 20.61 0.6064 - - 17.98 0.5269 - -
+Event-3DGS (E+F) E+F 23.43 0.6456 - - 20.04 0.5515 - -
+EvCF-3DGS E+F 22.89 0.6317 1.78 0.82 19.13 0.5380 2.70 1.28
+EF-3DGS(Ours) E+F 23.65 0.6466 1.41 0.69 21.12 0.5620 1.80 0.89
+Table 3: Rendering and pose estimation results on RealEv-DAVIS. Complete data and additional metrics are provided in the supplementary material.
+Results on RealEv-DAVIS.
+Table 3 validates our approach on the real-world RealEv-DAVIS dataset.
+EF-3DGS outperforms top-performing methods and handles real-world scenes effectively.
+In FAST scenarios, our method shows nearly 1dB PSNR improvement over the best baselines.
+This confirms our advantage in high-speed scenarios where frame-based methods struggle.
+Fig.
 4 and Fig.
-7 showcase EF-3DGS’s superior preservation of fine-grained details and the method’s robustness to real-world trajectory.
-V-D3 Performance under Varying Camera Speeds
-As evidenced in Table I to Table IV, the performance of all methods generally declines as the frame rate decreases.
-However, our method exhibits remarkable resilience to the reduction in frame rate, maintaining its superiority over the other methods across all metrics.
-Regarding novel view synthesis quality, the performance gap between the proposed method and the other methods consistently widens as the frame rate decreases.
-For example, the rendering quality difference between our method and CF-3DGS [5], measured in PSNR, increases from 0.61dB at 6FPS to 3.43dB at 1FPS.
-This observation demonstrates the robustness of EF-3DGS in handling challenging high-speed scenarios, where the advantages of leveraging event data become more pronounced.
-V-D4 Robustness to Pose Disturbance
-To further validate the performance of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP.
-Specifically, following BARF [42], we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra.
-For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level.
-Then, each method is initialized with the noised poses, after which the optimization is performed.
-The results are illustrated in Fig.
-11.
-Notably, ENeRF [43], which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases.
-This observation validates the critical importance of optimizing camera poses alongside the scene reconstruction.
-Furthermore, Our proposed framework exhibits superior tolerance across the entire range of pose perturbations compared to existing methods.
-Even in the presence of significant perturbation, the degradation in EF-3DGS’s novel view synthesis fidelity and trajectory estimation performance is substantially lower than that of the other methods.
-Figure 8: Qualitative Results of the effectiveness of $\mathcal{L}_{EGM}$.
-TABLE V: Effect of each component in EF-3DGS. Evaluation is conducted on 1FPS scenes of Tanks and Temples. The best results are highlighted in bold.
+5 show our method preserves fine details and maintains accurate trajectories even during rapid motion, addressing key limitations of traditional approaches.
+Performance under Varying Camera Speeds  As shown in Fig 2 and Fig.
+3, while all methods degrade as the frame rate decreases, our approach shows remarkable resilience.
+The performance gap widens significantly at lower frame rates, with our PSNR advantage over CF-3DGS [9] increasing from 0.61dB at 6FPS to 3.43dB at 1FPS.
+Notably, our method also consistently outperforms other event-based methods (EvCF-3DGS and Event-3DGS).
+This confirms not only the value of event data in challenging scenarios but also the superiority of our integration approach.
 $\mathcal{L}_{EGM}$ $\mathcal{L}_{LEGM}$ $\mathcal{L}_{PBA}$ Fixed GS NVS Pose
@@
 23.96 0.72 0.36 0.0487 0.6259 0.3753
-Figure 9: Qualitative Results of the effectiveness of $\mathcal{L}_{LEGM}$.
-Figure 10: Qualitative Results of the effectiveness of $\mathcal{L}_{PBA}$.
+Table 4: Effect of each component in EF-3DGS. The best results are highlighted in bold.
+5.5 Ablation Studies
+Effect of Each Component  Table 4 presents a comprehensive ablation study of our key components under the challenging 1FPS setting on Tanks and Temples.
+$\mathcal{L}_{EGM}$ serves as the foundation of our approach, providing substantial improvements in both rendering quality (+1.63dB PSNR) and pose accuracy by enabling rich supervision between discrete frames.
+Building upon this, $\mathcal{L}_{LEGM}$ extracts motion information from events and constrains 3DGS in the gradient domain, significantly improving pose estimation while modestly enhancing rendering quality.
+$\mathcal{L}_{PBA}$, though designed to address color inconsistency issues, not only improves rendering quality but also enhances pose estimation accuracy by establishing geometric and photometric consistency across views.
+The Fixed-GS training strategy, while having no impact on pose optimization, significantly improves rendering quality by effectively separating structure and color optimization.
+We provide more intuitive ablation visualizations in the supplementary material.
 Figure 6: Robustness of different methods to pose disturbance.
-TABLE VI: Effect of the number of event subintervals N.
-[…]
-$N=2$ 23.56 30+ 1.3h
-$N=3$ 24.43 30+ 3h
-$N=6$ 24.81 30+ 7h
-[…]
-CF-3DGS 22.08 30+ 3h
-TABLE VII: Comparison of training time vs PSNR for various methods
-V-E Ablation Studies
-To further investigate the effectiveness of each key component in our proposed EF-3DGS framework, we conduct ablation experiments under the 1 FPS setting on the Tanks and Temples dataset.
-The results are presented in Table V.
-V-E1 Effect of $\mathcal{L}_{EGM}$
-Analysis of Table V demonstrates that removing all components substantially degrades both rendering quality and pose estimation accuracy.
-This finding underscores the significant challenges faced by the original CF-3DGS in high-speed scenarios.
-Fig.
-8 further illustrates this issue, revealing that while CF-3DGS can produce high-quality renderings at training views, it struggles with adjacent novel views.
-Moreover, the rendered depth maps expose considerable inaccuracies in CF-3DGS’s scene geometry reconstruction.
-The incorporation of $\mathbf{\mathcal{L}_{EGM}}$ yields marked improvements in both rendering quality and pose estimation precision, as evidenced by Table V and Fig.
-8.
-This experimental outcome suggests that $\mathbf{\mathcal{L}_{EGM}}$ effectively harnesses the continuous brightness change data captured by event cameras, thereby compensating for the inter-frame loss of scene brightness information.
-Consequently, the 3DGS training optimization achieves multi-view geometric consistency.
-Notably, among all components, $\mathbf{\mathcal{L}_{EGM}}$ demonstrates the most profound impact on overall method performance, establishing its role as a fundamental component of our approach.
-V-E2 Effect of $\mathcal{L}_{LEGM}$
-As demonstrated in Table V, $\mathbf{\mathcal{L}_{LEGM}}$ further enhances both rendering quality and pose estimation accuracy when applied in conjunction with $\mathbf{\mathcal{L}_{EGM}}$.
-Fig.
-9 provides a visual illustration of $\mathbf{\mathcal{L}_{LEGM}}$’s efficacy.
-Ideally, with accurate depth and pose estimation, the Image of Warped Events (IWE) should resemble the reference image.
-However, in the absence of $\mathbf{\mathcal{L}_{LEGM}}$, the generated IWE exhibits blurring effects, indicative of inaccurate pose estimation.
-More critically, as highlighted in the boxed area of Fig.
-9, the presence of self-similar regions with repetitive textures poses additional challenges for scene reconstruction and pose estimation.
-Significant errors in scene reconstruction leading to inaccurate depth estimates further compromise the inherent structure of the IWE.
-The introduction of $\mathbf{\mathcal{L}_{LEGM}}$ allows for the simultaneous reflection of both pose estimation and scene reconstruction errors in the IWE.
-By optimizing the sharpness of the IWE, $\mathbf{\mathcal{L}_{LEGM}}$ effectively improves both rendering quality and pose estimation precision.
-V-E3 Effect of $\mathcal{L}_{PBA}$ and Fixed-GS Strategy
-Comparing the last three rows of Table V, we observe that both $\mathbf{\mathcal{L}_{PBA}}$ and Fixed-GS strategies significantly enhance rendering quality.
-$\mathbf{\mathcal{L}_{PBA}}$, by enforcing multi-view consistency, slightly improves pose estimation accuracy, while the Fixed-GS strategy, applied solely during scene optimization, has no impact on pose estimation.
-Fig.
-10 visualizes the effects of these components.
-When both are removed, the reconstructed scene loses almost all color information.
-This is attributed to the substantially higher number of event frames than image frames, causing event-guided optimization to dominate scene reconstruction and suppress color information from image frames.
-$\mathbf{\mathcal{L}_{PBA}}$ alone, which reprojects neighboring video frames onto event frames, only partially mitigates this imbalance.
-The Fixed-GS strategy employs a two-stage scene optimization approach for the separation of the color and structure of the scene.
-However, this approach does not fully leverage the supervisory signals from adjacent frames.
-The combination of $\mathbf{\mathcal{L}_{PBA}}$ and Fixed-GS strategies complements each other, more effectively addressing the color distortion issue.
-V-E4 Computational Efficiency and Impact of Subinterval Number N
-To investigate the impact of the number of subintervals N on our method, we conduct ablation studies on the Tanks and Temples dataset under the 2 FPS setting.
-The results are shown in Table VII and figure on the right.
-We test the speed on RTX2080ti.
-Note that in the t-PSNR figure, circular nodes represent pose-free methods, while diamond-shaped nodes indicate methods that require precomputed poses.
-The results in Table VII reveal that increasing N improves PSNR, indicating that finer temporal discretization enhances reconstruction quality.
-However, this improvement comes at the cost of extended training time, increasing from 1.3 hours at N=2 to 7 hours at N=6.
-Notably, our method maintains real-time rendering performance, thanks to the efficient rendering capabilities of 3DGS.
-Even at N=2, our approach outperforms baselines in both PSNR and rendering speed.
-The setting of N=3 offers a good balance, achieving higher PSNR than CF-3DGS with comparable training time.
-However, it’s important to note that the optimal N may vary for different frame rates and motion characteristics.
-Figure 12: Study of the different evalution metrics with respect to $\lambda_{cm}$
-Figure 13: Study of the different evalution metrics with respect to $\lambda_{grad}$
-Figure 14: Study of the different evalution metrics with respect to $\lambda_{PBA}$
-V-F Analysis of Loss Coefficients
-Different components of our method play crucial roles in overall performance.
-To comprehensively understand these effects and determine optimal weight settings, we conducted a detailed analysis of the loss weights: $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$.
-V-F1 Contrast Maximization Coefficient $\lambda_{cm}$
-As shown in Fig.
-14, the contrast maximization coefficient $\lambda_{cm}$ significantly affects both NVS quality and pose estimation.
-We observe that a value of 0.1 achieves the best overall performance.
-Lower values (0.01) lead to decreased performance, likely due to insufficient utilization of motion information from events.
-Higher values (0.5) also result in performance degradation, possibly due to over-reliance on the contrast maximization term at the expense of other constraints.
-The optimal $\lambda_{cm}$ suggests that while event data is crucial, it should not dominate the reconstruction process.
-This balance allows our method to leverage the high temporal resolution of events without sacrificing the global supervision provided by frame data.
-V-F2 Gradient-based Loss Coefficient $\lambda_{grad}$
-As shown in Fig.
-14, the gradient-based loss coefficient $\lambda_{grad}$ shows optimal performance at 0.2.
-Lower values (0.1) slightly decrease performance, while higher values (0.3, 0.5) lead to more significant drops in both NVS quality and pose accuracy.
-This is likely because $\mathcal{L}_{EGM}$ and 3DGS primarily focus on rendering absolute brightness, and excessive gradient constraints may interfere with this process.
-Therefore, a moderate $\lambda_{grad}$ value is crucial to balance gradient information with the primary rendering objectives.
-V-F3 Photometric Bundle Adjustment Coefficient $\lambda_{PBA}$
-As illustrated in Figure 14, our method demonstrates remarkable stability across $\lambda_{PBA}$ values ranging from 0.1 to 0.5.
-However, Fig.
-10 illustrates that while $\mathcal{L}_{PBA}$ has minimal effect on PSNR and ATE metrics, its absence leads to noticeable color distortion artifacts in visual results.
-This underscores the importance of $\lambda_{PBA}$ in maintaining visual fidelity.
-Conversely, setting $\lambda_{PBA}$ to 1.0 results in significant performance degradation, indicating that overemphasis on cross-view consistency can be counterproductive.
-VI Limitations
-Our method combines event streams with conventional video frames, which require the input data to be time-ordered.
-So it can not handle the unordered image input.
-In high-speed situations, video frames may suffer from severe motion blur, which can degrade the quality of the visual information.
-Future research could focus on this challenge.
-And, our method relies on several key parameters (e.g., $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$) that may require manual tuning for optimal performance across different speed scenarios.
-This dependency on scene-specific parameter settings could limit the method’s adaptability to diverse environments.
-Future work should explore self-adaptive parameter adjustment strategies to enhance the method’s versatility and ease of use across various reconstruction tasks.
-VII Conclusions
+Robustness to Pose Disturbance  To validate the robustness of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP.
+Specifically, following BARF [23], we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra.
+For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level.
+Then, each method is initialized with the noised poses, after which the optimization is performed.
+The results are illustrated in Fig.
+6.
+Notably, Event-3DGS [16], which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases.
+This observation validates the critical importance of joint pose-scene optimization.
+Furthermore, Our proposed framework demonstrates superior tolerance across all perturbation levels.
+Even under significant noise, our method experiences substantially less degradation in both rendering quality and trajectory accuracy.
+6 Conclusions
 In this work, we propose Event-Aided Free-Trajectory 3DGS (EF-3DGS), a novel framework that seamlessly integrates event camera data into the task of reconstructing 3DGS from casually captured free-trajectory videos.

### v4

@@
 Figure 1: Free-trajectory 3DGS under high speed.
-The top row shows the overall paradigm.
-The colored dots (red for positive and blue for negative events) in the top row represent the event data.
+(Top) The overall paradigm.
+The colored dots in the top row represent the event data (red: positive, blue: negative).
 We leverage continuous event streams to aid discrete video frames captured along free trajectories in high-speed scenarios, jointly optimizing camera poses and reconstructing the 3DGS.
-Our method surpasses current state-of-the-art methods in terms of both rendered results (middle row) and pose estimation (bottom row).
+Our method surpasses current state-of-the-art methods in terms of both rendered results (middle) and pose estimation (bottom).
 1 Introduction
-In recent years, Neural Radiance Fields (NeRF) [30, 1, 31] and 3D Gaussian splatting (3DGS) [20, 9] have made significant progress in novel view synthesis tasks.
+In recent years, Neural Radiance Fields (NeRF) nerf ; mipnerf360 ; instantngp and 3D Gaussian splatting (3DGS) 3dgs ; cf3dgs have made significant progress in novel view synthesis tasks.
 Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering.
-While subsequent methods [1, 31, 7, 2] excel with posed images, reconstructing scenes from videos with free camera trajectories remains challenging despite its applications in VR/AR, video stabilization, and mapping.
+While subsequent methods mipnerf360 ; instantngp ; tensorf ; zipnerf excel with posed images, reconstructing scenes from videos with free camera trajectories remains challenging despite its applications in VR/AR, video stabilization, and mapping.
 To tackle this challenging task, several efforts have been made.
@@
 One line of work draws inspiration from Simultaneous Localization and Mapping (SLAM).
-They [28, 9, 19] follow its optimization paradigm, progressively optimizing camera trajectories and alternating between camera pose and scene refinement.
-Another line of work [4, 19, 25, 9, 42] explores incorporating additional geometric or motion priors such as depth estimation [29, 36] or optical flow [41] to establish constraints beyond photometric rendering loss.
+They localrf ; cf3dgs ; cogs follow its optimization paradigm, progressively optimizing camera trajectories and alternating between camera pose and scene refinement.
+Another line of work nopenerf ; cogs ; rodynrf ; cf3dgs ; sparf explores incorporating additional geometric or motion priors such as depth estimation depth1 ; depth2 or optical flow raft to establish constraints beyond photometric rendering loss.
 While these methods can render photo-realistic images in typical free-trajectory scenarios, both their rendering quality and pose estimation accuracy degrade significantly in high-speed scenarios (or equivalently low-frame-rate scenarios) as shown in Fig.
@@
 First, the limited number of camera observations leads to an under-constrained scene reconstruction problem.
-This can cause the scene representation to converge to a trivial solution [32, 47, 42], where the model overfits to the training views without capturing the correct underlying geometry structure.
+This can cause the scene representation to converge to a trivial solution regnerf ; sparsenerf ; sparf , where the model overfits to the training views without capturing the correct underlying geometry structure.
 Second, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which is leveraged by previous methods.
@@
 These significant violations greatly exacerbate the ill-posedness of the joint optimization of scene and camera poses.
-Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur [12, 17, 45, 46, 52, 6].
+Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur event_survey ; EDSO ; ultimateslam ; wan2024event ; wu2024event ; chengzhi .
 The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames.
-Moreover, the event data naturally encodes the motion information of the scene [24, 40, 44], containing rich motion cues.
+Moreover, the event data naturally encodes the motion information of the scene EvMotionEstimation ; ganchao ; 10047993 , containing rich motion cues.
 These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios.
@@
 This allows us to utilize the brightness cues encoded in the event stream between adjacent frames, providing rich supervisory signals to alleviate the insufficient sparse view issues.
-(2) In the Linear Event Generation Model (LEGM), regarding the pose estimation challenge, we introduce the CMax [11] framework to exploit the spatio-temporal correlations of events.
+(2) In the Linear Event Generation Model (LEGM), regarding the pose estimation challenge, we introduce the CMax cmax framework to exploit the spatio-temporal correlations of events.
 We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive frames.
 We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns.
-Furthermore, through the Linear Event Generation Model [10, 13], we establish a connection between motion and brightness changes.
+Furthermore, through the LEGM legm2015 ; legm2018 , we establish a connection between motion and brightness changes.
 This allows us to constrain the 3DGS in the gradient domain using the IWE.
@@
 2 Related Works
-2.1 Joint Pose and Scene Optimization
-The research community has recently focused on developing methods [42, 28, 50, 8, 4, 23, 9] that can be optimized without requiring precomputed camera poses.
+Joint Pose and Scene Optimization.
+The research community has recently focused on developing methods sparf ; localrf ; nerfmm ; graf ; nopenerf ; barf ; cf3dgs that can be optimized without requiring precomputed camera poses.
 A line of work has focused on improving the stability of the optimization process.
-GARF [8] and BARF [23] both find that the high-frequency position encoding is prone to local minima and try to improve it.
-For example, GARF [8] proposes using Gaussian activation to replace the sinusoidal position encoding.
+GARF graf and BARF barf both find that the high-frequency position encoding is prone to local minima and try to improve it.
+For example, GARF graf proposes using Gaussian activation to replace the sinusoidal position encoding.
 Another line of work has investigated incorporating additional constraints to make the problem more tractable.
-LocalRF [28] leverages the prior assumption of continuous motion between adjacent frames and progressively adds and optimizes camera poses.
-More recent approaches [4, 28, 9] leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation.
-Exploiting 3DGS’s explicit representation, CF-3DGS [9] directly back-projects Gaussian points using depth maps.
+LocalRF localrf leverages the prior assumption of continuous motion between adjacent frames and progressively adds and optimizes camera poses.
+More recent approaches nopenerf ; localrf ; cf3dgs leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation.
+Exploiting 3DGS’s explicit representation, CF-3DGS cf3dgs directly back-projects Gaussian points using depth maps.
 While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed scenarios or rely on a good pose initialization.
 Our approach addresses these issues by leveraging motion and brightness cues from event streams.
-2.2 Event-Based Novel View Synthesis
-Recent works have explored the integration of event cameras into the NeRF or 3DGS framework.
-Early approaches, such as E-NeRF [21] and EventNeRF [37], utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes.
-Building upon this, Robust e-NeRF [26] incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise.
+Event-Based Novel View Synthesis.
+Recent works have explored the integration of event cameras EMoTive ; etap ; urnet into the NeRF or 3DGS framework.
+Early approaches, such as E-NeRF enerf and EventNeRF eventnerf , utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes.
+Building upon this, Robust e-NeRF robustenerf incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise.
 Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods.
-For instance, E2NeRF [35] and EvDeblurNeRF [5] leverage the Event Double Integral (EDI) [33] model to address the deblurring problem, while DE-NeRF [27] and EvDNeRF [3] leverage the high temporal resolution property of event cameras to capture fast-moving elements in dynamic scene.
-More recently, Event-3DGS [16] and EaDeblur-GS [51] have extended previous approaches to 3D Gaussian Splatting, achieving superior rendering quality and real-time performance.
+For instance, E2NeRF E2NeRF and EvDeblurNeRF evdeblurnerf leverage the Event Double Integral (EDI) edi model to address the deblurring problem, while DE-NeRF denerf and EvDNeRF EvDNeRF leverage the high temporal resolution property of event cameras to capture fast-moving elements in dynamic scene.
+More recently, Event-3DGS event3dgs and EaDeblur-GS eadeblurgs have extended previous approaches to 3D Gaussian Splatting, achieving superior rendering quality and real-time performance.
 A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we target free-trajectory scenarios, jointly optimizing for both the camera poses and the scene representation.
@@
 Figure 2: Method overview.
-Our method takes video frames and event stream as input.
-During the first stage, we progressively add new event images, leveraging the events and most recent frame to establish the event-driven optimization.
-During the second stage, we adopt the Fixed-GS strategy to mitigate the color distortion of 3DGS.
-The details of $\mathcal{L}_{LEGM}$ and CMax framework are illustrated in Fig.
+The inputs are video frames and event stream.
+In the first stage, we progressively add new event images, leveraging the events and most recent frame to establish the event-driven optimization.
+In the second stage, we adopt the Fixed-GS strategy to mitigate the color distortion of 3DGS.
+The details of $\mathcal{L}_{LEGM}$ and CMax framework are shown in Fig.
 3.
 3 Preliminary
-3DGS [20] parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information.
+3DGS 3dgs parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information.
 Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$.
@@
 where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel.
-With the forward rendering procedure, we can optimize 3DGS by minimizing a weighted combination loss of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ between observation and rendered pixels: $\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ where $\lambda$ is balancing weight which is set to 0.2 following [20].
-By integrating depth $d_{i}$ in Equation 1 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
+With the forward rendering procedure, we can optimize 3DGS by minimizing a weighted combination loss of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ between observation and rendered pixels: $\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ where $\lambda$ is balancing weight which is set to 0.2 following 3dgs .
+By integrating depth $d_{i}$ in Equation˜1 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
 $\hat{D}(\mathbf{r})=\sum_{i=1}^{N}d_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}).$ (2)
 4 Method
-The overall framework is shown in Fig.
-2.
-Given a video of a free-trajectory $\{I_{i}\}$ captured at time $\{t_{i}\}$ and the event stream $\varepsilon=\{\mathbf{e}_{k}\}$, our goal is to reconstruct the 3DGS of the scene and the corresponding camera trajectory$\{T_{i}\}$.
-Following the analysis-by-synthesis paradigm of 3DGS, we extend this approach by incorporating event camera data through two fundamental imaging principles: the Event Generative Model (EGM) and Linear Event Generative Model (LEGM).
-To address the absence of color information in events and ensure cross-view consistency, we further introduce photometric bundle adjustment (PBA) and a Fixed-GS training strategy.
+[…]
 4.1 EGM Driven Optimization
-The Event Generation Model (EGM) describes how event cameras asynchronously record pixel-wise brightness changes.
+The EGM describes how event cameras asynchronously record pixel-wise brightness changes.
 When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,
@@
 By enforcing consistency between rendered and latent intensity images, this loss effectively utilizes the brightness information encoded in event streams between adjacent frames, addressing the challenge of sparse viewpoints in high-speed scenarios.
+Figure 3: The illustration of unified CMax and LEGM optimization.
+We warp previous event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of IPWE.
+The byproduct IPWE is utilized to establish additional constraints on 3DGS.
 4.2 Unified CMax and LEGM Optimization
 While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by events, it does not explicitly exploit the motion information encoded in the event stream.
-To address this, we introduce the Contrast Maximization (CMax) [11, 34, 38, 15] framework and the Linear Event Generation Model (LEGM) [10, 13, 56].
+To address this, we introduce the Contrast Maximization (CMax) cmax ; cmax2 ; cmax3 ; cmaxslam framework and the LEGM legm2015 ; legm2018 ; linearinversed .
 These models complement the previous EGM-driven optimization.
@@
 Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the estimated motion.
-This insight motivates us to derive the motion field by leveraging the rendered depth from 3DGS using Eq.
-2 and the relative camera motion between neighboring timestamps.
+This insight motivates us to derive the motion field by leveraging the rendered depth from 3DGS using eq.˜2 and the relative camera motion between neighboring timestamps.
 By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the camera poses.
@@
 Specifically, for current timestamp $t_{ref}=t_{i,j}$, we warp the event frames from previous $r$ sub-intervals:
-$\vspace{1pt}E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),\vspace{1pt}$ (7)
-where $m\in[0,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow derived from the rendered depth $\hat{D}$ in  Eq. 2 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
-$\vspace{1pt}F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),\vspace{1pt}$ (8)
+$E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),$ (7)
+where $m\in[0,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow derived from the rendered depth $\hat{D}$ in  eq.˜2 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
+$F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),$ (8)
 where $T_{i,j\rightarrow j-m}=T_{i,j-m}T_{i,j}^{-1}$, $\Pi$ projects a 3D point to image coordinates and $\Pi^{-1}$ unprojects a pixel coordinate and depth into a 3D point.
@@
 $\mathcal{L}_{cm}=-\textrm{Var}(\textrm{IPWE}_{i,j}).$ (10)
-Furthermore, based on the Linear Event Generation Model (LEGM) [10, 13], the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of an image):
+Furthermore, based on the LEGM legm2015 ; legm2018 , the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of an image):
 $\Delta L(\mathbf{u})=-\nabla L\cdot\dot{\mathbf{u}}\approx L(\mathbf{u})-L(\mathbf{u}+\dot{\mathbf{u}}).$ (11)
-It is noteworthy that the IPWE also encodes brightness change information. Combining  Eq. 9 and  Eq. 11, we establish a connection between the IPWE and the brightness changes of the rendered images:
+It is noteworthy that the IPWE also encodes brightness change information. Combining  eq.˜9 and  eq.˜11, we establish a connection between the IPWE and the brightness changes of the rendered images:
 $C\cdot\textrm{IPWE}_{i,j}=\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}).$ (12)
-Note that to compute $F_{i,j\rightarrow j+1}$, we estimate $T_{i,j,j+1}$ by leveraging the assumption of locally linear motion from $T_{i,j,j-1}$ and $T_{i,j,j}$.
+Note that to compute $F_{i,j\rightarrow j+1}$, we estimate $T_{i,j+1}$ by leveraging the assumption of locally linear motion from $T_{i,j-1}$ and $T_{i,j}$.
 Based on this relationship, we formulate an additional gradient-based loss:
-$\vspace{1pt}\mathcal{L}_{grad}=||C\cdot\textrm{IPWE}_{i,j}-(\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}))||^{2},\vspace{1pt}$ (13)
+[…]
 where $\hat{L}$ is the logarithm of synthesised image $\hat{I}_{t}$. Finally, the full LEGM loss is defined as:
@@
 Specifically, as shown in Fig. 2 (1.3), for a randomly sampled timestamp $t\in\{t_{i,j}\}$, we establish the following photometric reprojection error:
-$\mathcal{L}_{PBA}=\sum_{\mathbf{u}\in\mathcal{P}}\sum_{I_{s}\in\mathcal{F}}||I_{s}(\mathbf{u}^{\prime})-\hat{I}(\mathbf{u})||^{2},$ (15)
+[…]
 where $\mathbf{u}^{\prime}=\Pi(T_{i,j-r\rightarrow j}\Pi^{-1}(x,y,\hat{D}(u)))$ represent the coordinate on target view projected from the pixel $\mathbf{u}$ of source view $I_{s}$, $\mathcal{P}$ denotes the pixel samples of current frame, and $\mathcal{F}$ is the candidates of target video frames.
@@
 By minimizing $\mathcal{L}_{PBA}$ across sampled views, we encourage the 3DGS model to produce geometrically and photometrically consistent renderings across events and video frames, thus effectively resolving color inconsistencies inherent in event data.
-Figure 3: The illustration of unified CMax and LEGM optimization.
-We warp previous event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of piece-wise warped events (IPWE).
-The byproduct IPWE is utilized to establish additional constraints on 3DGS.
+4.4 Fixed-GS Training Strategy
+The $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues.
+To further address this challenge, we propose a two-stage Fixed-GS scene optimization strategy that takes advantage of 3DGS’s explicit attribute representation.
+In the first stage, all the parameters are optimizable and the optimization is performed across all timestamps:
+$G_{\theta}^{*},T_{i,j}^{*}=\underset{\mu,\alpha,r,s,f,T_{i,j}}{\operatorname{argmin}}~\mathcal{L}_{event},~t\in\{t_{i,j}\},$ (16)
+where $\mu,\alpha,r,s,f$ is the position, opacity, rotation, scale factor and spherical harmonics of the Gaussians, and t is the sampled timestamp during training.
+This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions due to the dominant colorless event supervision overwhelming the sparse RGB frame color supervision.
+The second stage focuses on recovering accurate color information.
+During this phase, optimization is conducted exclusively on video frames.
+We optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed:
+$G_{\theta}^{*}=\underset{f}{\operatorname{argmin}}~\mathcal{L}_{color},~t\in\{t_{i,0}\}$ (17)
+The ratio between the first and second stages is empirically set to 4:1.
+This approach allows us to effectively address the color distortion problem while preserving the structural and brightness information obtained from the event data.
+4.5 Overall Training Pipeline
+Assembling all loss terms, we get the overall loss function:
+$\mathcal{L}_{event}=\mathcal{L}_{EGM}+\mathcal{L}_{LEGM}+\lambda_{PBA}\mathcal{L}_{PBA},$ (18)
+where $\lambda_{PBA}$ are the weighting factor.
+Note that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space.
+We incorporate dynamic scene allocation strategies from LocalRF localrf for handling extended video sequences.
+Our overall training pipeline builds upon the progressive optimization scheme of CF-3DGS cf3dgs while introducing novel components to integrate event stream data for robust free-trajectory scene reconstruction.
+Please refer to Section˜A.3 for the algorithm pipeline and additional implementation details.
+5 Experiments
 Methods Pose-Free Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
@@
 […]
-Table 1: Quantitative evaluations on Tanks and Temples.
-Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol.
+Table 1: Quantitative evaluations on Tanks and Temples dataset.
 The best results are highlighted in bold.
+Note that, in the “input” column, “F” denotes traditional frame input, while “E+F” denotes hybrid frame and event input.
 Methods Input 6 FPS 4 FPS 3 FPS 2 FPS 1 FPS
@@
 Those methods that require precomputed poses are excluded.
-4.4 Fixed-GS Training Strategy
-The $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues.
-To further address this challenge, we propose a two-stage Fixed-GS scene optimization strategy that takes advantage of 3DGS’s explicit attribute representation.
-In the first stage, all the parameters are optimizable and the optimization is performed across all timestamps:
-$G_{\theta}^{*},T_{i,j}^{*}=\underset{\mu,\alpha,r,s,f,T_{i,j}}{\operatorname{argmin}}~{}\mathcal{L}_{event},~{}t\in\{t_{i,j}\},$ (16)
-where $\mu,\alpha,r,s,f$ is the position, opacity, rotation, scale factor and spherical harmonics of the Gaussians, and t is the sampled timestamp during training.
-This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions due to the dominant colorless event supervision overwhelming the sparse RGB frame color supervision.
-The second stage focuses on recovering accurate color information.
-During this phase, optimization is conducted exclusively on video frames.
-We optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed:
-[…]
-The ratio between the first and second stages is empirically set to 4:1.
-This approach allows us to effectively address the color distortion problem while preserving the structural and brightness information obtained from the event data.
-4.5 Overall Training Pipeline
-Assembling all loss terms, we get the overall loss function:
-$\mathcal{L}_{event}=\mathcal{L}_{EGM}+\mathcal{L}_{LEGM}+\lambda_{PBA}\mathcal{L}_{PBA},$ (18)
-where $\lambda_{PBA}$ are the weighting factor.
-Note that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space.
-We incorporate dynamic scene allocation strategies from LocalRF [28] for handling extended video sequences.
-Our overall training pipeline builds upon the progressive optimization scheme of CF-3DGS [9] while introducing novel components to integrate event stream data for robust free-trajectory scene reconstruction.
-Please refer to the Supplementary Material for the algorithm pipeline and additional implementation details.
-5 Experiments
-Figure 4: Qualitative comparison for novel view synthesis.
-The first two rows are from Tanks and Temples and the last row is from RealEv-DAVIS.
-Our approach produces more realistic rendering results with fine-grained details.
-Better viewed when zoomed in.
-Figure 5: Pose estimation comparison.
-We visualise the trajectory (3D plot) and $\textrm{RPE}_{r}$ (color bar) of each method.
-We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
 5.1 Dataset
 Tanks and Temples.
-We conduct comprehensive experiments on the Tanks and Temples dataset [22].
-Similar to LocalRF [28], we adopt 9 scenes, covering large-scale indoor and outdoor scenes.
+We conduct comprehensive experiments on the Tanks and Temples dataset tnt .
+Similar to LocalRF localrf , we adopt 9 scenes, covering large-scale indoor and outdoor scenes.
 For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance.
-Following LocalRF [28], we apply 4× spatial downsampling to the videos.
+Following LocalRF localrf , we apply 4× spatial downsampling to the videos.
 To evaluate the robustness under varying camera speeds, we employ varying temporal downsampling of 6 FPS, 4 FPS, 3 FPS, 2 FPS, and 1 FPS.
 The reduction in frame rate effectively creates larger inter-frame displacements, simulating high-speed scenarios.
-To synthesize realistic event data, we first upsample the original videos by [18] and then apply the simulator V2E [14].
+To synthesize realistic event data, we first upsample the original videos by RIFE and then apply the simulator V2E v2e .
 RealEv-DAVIS.
-Due to the lack of free-trajectory event camera datasets, we introduce RealEv-DAVIS, comprising four outdoor scenes.
+Due to the lack of free-trajectory event camera datasets, we introduce RealEv-DAVIS.
 Using a DAVIS346 camera that simultaneously captures frames and events at 346×260 resolution, we record 40-second handheld sequences at 25 FPS.
@@
 For SLOW scenarios, we retain every second frame, while for FAST scenarios, we keep only one frame per five frames.
-Further details are provided in the supplementary material.
+Further details are provided in Section˜A.2.
 5.2 Implementation details
-We follow the optimization parameters by the configuration outlined in the 3DGS [20].
+We follow the optimization parameters by the configuration outlined in the 3DGS 3dgs .
 We optimize the camera poses in the representation of quaternion rotation.
@@
 The contrast threshold $C$ is set to 0.25 for Tanks and Temples and 0.21 for RealEv-DAVIS.
-We provide detailed ablation studies on these hyperparameters and additional implementation details in the supplementary material.
+We provide detailed ablation studies on these hyperparameters and additional implementation details in Section˜A.4.
+Figure 4: Qualitative comparison for novel view synthesis.
+The first two rows are from Tanks and Temples and the last row is from RealEv-DAVIS.
+Our approach produces more realistic rendering results with fine-grained details.
+Better viewed when zoomed in.
+Figure 5: Pose estimation comparison.
+We visualise the trajectory (3D plot) and $\mathrm{RPE}_{r}$ (color bar) of each method.
+We clip and normalize the $\mathrm{RPE}_{r}$ by a quarter of the max $\mathrm{RPE}_{r}$ across all results of each scene.
 5.3 Experimental Setup
@@
 We evaluate all the methods from two aspects: novel view synthesis and pose estimation.
-For the novel view synthesis task, we report the standard metrics PSNR, SSIM [49], and LPIPS [54].
-For the pose estimation task, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [55, 39], as delineated in [4].
+For the novel view synthesis task, we report the standard metrics PSNR, SSIM ssim , and LPIPS lpips .
+For the pose estimation task, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics pose_eval1 ; pose_eval2 , as delineated in nopenerf .
 Since these metrics are inherently influenced by frame rate, we upsample all estimated poses to a consistent temporal resolution before evaluation for fair comparison across different frame rate settings.
 Baselines.
-For a fair comparison, we focus on two categories of methods: (1) For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF [28] and F2-NeRF [48].
-We also include pose-free methods like Nope-NeRF [4] and CF-3DGS [9].
-(2) For event-frame hybrid methods, we consider approaches that fuse events and frames, including ENeRF [21], EvDeblurNeRF [5] and Event-3DGS [16].
-Since no existing method integrates events for free-trajectory scenarios, we implement EvCF-3DGS as a competitive baseline that leverages an event-based frame interpolation network (Time Lens [43]) to temporally upsample frames before feeding them into CF-3DGS.
-5.4 Experimental Results
-We select every ten frames as a test image for NVS evaluation following LocalRF [28].
-Since the camera poses are unknown in our setting, we need to estimate the poses of test views.
-As in iNeRF [53], we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
-Results on Tanks and Temples.
-Tables 1 and 2 demonstrate two key findings: (1) Our event-aided approach achieves up to 3dB higher PSNR and nearly 40% lower trajectory error at 1FPS compared to frame-based methods, indicating the critical value of event data in high-speed scenarios.
-(2) Our method maintains 1.55dB PSNR advantage over Event-3DGS at 1FPS, confirming that our integration framework effectively exploits the nature of event data beyond merely using it.
-Fig.
-4 shows our method produces sharper edges and finer textures, while Fig.
-5 illustrates we achieve more accurate trajectory estimation.
+For a fair comparison, we focus on two categories of methods: (1) For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF localrf and F2-NeRF f2nerf .
+We also include pose-free methods like Nope-NeRF nopenerf and CF-3DGS cf3dgs .
+(2) For event-frame hybrid methods, we consider approaches that fuse events and frames, including ENeRF enerf , EvDeblurNeRF evdeblurnerf and Event-3DGS event3dgs .
 Methods Input SLOW FAST
@@
 Table 3: Rendering and pose estimation results on RealEv-DAVIS. Complete data and additional metrics are provided in the supplementary material.
+Since no existing method integrates events for free-trajectory scenarios, we implement EvCF-3DGS as a competitive baseline that leverages an event-based frame interpolation network (Time Lens timelens ) to temporally upsample frames before feeding them into CF-3DGS.
+5.4 Experimental Results
+We select every ten frames as a test image for NVS evaluation following LocalRF localrf .
+Since the camera poses are unknown in our setting, we need to estimate the poses of test views.
+As in iNeRF iNeRF , we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
 Results on RealEv-DAVIS.
@@
 5 show our method preserves fine details and maintains accurate trajectories even during rapid motion, addressing key limitations of traditional approaches.
-Performance under Varying Camera Speeds  As shown in Fig 2 and Fig.
-3, while all methods degrade as the frame rate decreases, our approach shows remarkable resilience.
-The performance gap widens significantly at lower frame rates, with our PSNR advantage over CF-3DGS [9] increasing from 0.61dB at 6FPS to 3.43dB at 1FPS.
-Notably, our method also consistently outperforms other event-based methods (EvCF-3DGS and Event-3DGS).
-This confirms not only the value of event data in challenging scenarios but also the superiority of our integration approach.
 $\mathcal{L}_{EGM}$ $\mathcal{L}_{LEGM}$ $\mathcal{L}_{PBA}$ Fixed GS NVS Pose
@@
 Table 4: Effect of each component in EF-3DGS. The best results are highlighted in bold.
+Figure 6: Robustness of different methods to pose disturbance.
+Results on Tanks and Temples.
+Tables 1 and 2 demonstrate two key findings: (1) Our event-aided approach achieves up to 3dB higher PSNR and nearly 40% lower trajectory error at 1FPS compared to frame-based methods, indicating the critical value of event data in high-speed scenarios.
+(2) Our method maintains 1.55dB PSNR advantage over Event-3DGS at 1FPS, confirming that our integration framework effectively exploits the nature of event data beyond merely using it.
+Fig.
+4 shows our method produces sharper edges and finer textures, while Fig.
+5 illustrates we achieve more accurate trajectory estimation.
+Performance under Varying Camera Speeds  As shown in Table 2 and Table 3, while all methods degrade as the frame rate decreases, our approach shows remarkable resilience.
+The performance gap widens significantly at lower frame rates, with our PSNR advantage over CF-3DGS cf3dgs increasing from 0.61dB at 6FPS to 3.43dB at 1FPS.
+Notably, our method also consistently outperforms other event-based methods (EvCF-3DGS and Event-3DGS).
+This confirms not only the value of event data in challenging scenarios but also the superiority of our integration approach.
 5.5 Ablation Studies
@@
 The Fixed-GS training strategy, while having no impact on pose optimization, significantly improves rendering quality by effectively separating structure and color optimization.
-We provide more intuitive ablation visualizations in the supplementary material.
-Figure 6: Robustness of different methods to pose disturbance.
+We provide more intuitive ablation visualizations in Section˜A.5.
 Robustness to Pose Disturbance  To validate the robustness of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP.
-Specifically, following BARF [23], we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra.
+Specifically, following BARF barf , we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra.
 For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level.
@@
 6.
-Notably, Event-3DGS [16], which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases.
+Notably, Event-3DGS event3dgs , which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases.
 This observation validates the critical importance of joint pose-scene optimization.
@@
 Our method effectively leverages the high temporal resolution and motion information encoded in event streams to enhance the 3DGS optimization process, leading to improved rendering quality and accurate camera pose estimation.
+By introducing the Event Generation Model and Linear Event Generation Model, we bridge the gap between differential event data and absolute brightness rendering in 3DGS.
+The proposed photometric bundle adjustment and Fixed-GS strategy further ensure accurate color recovery and scene structure optimization.
+Extensive experiments on both public benchmarks and real-world datasets validate the effectiveness of our approach, demonstrating significant improvements over state-of-the-art methods in both rendering quality and trajectory estimation accuracy.
+Future work would explore self-adaptive parameter adjustment strategies to enhance the method’s versatility and ease of use across various reconstruction tasks.
+7 Acknowledgements
+This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62225207, 62436008 and 62306295.
+The AI-driven experiments, simulations and model training were performed on the robotic AI-Scientist platform of Chinese Academy of Sciences.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The abstract and introduction accurately reflect the main contributions. Please refer to the main body of the paper.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Justification: Please refer to Appendix A.9 for our discussions on the limitations.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We provide the necessary formulas and related explanations in the main text. Please see refer to Method section about the blur model in main paper.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Justification: We identify our baseline method and explain the implementation details where we differ from the baseline method in Experiment Setup 5.3.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24. Justification: The datasets and baseline methods in this paper are publicly available. Our source code and the self-collect dataset will be openly accessible after the paper is accepted.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We have provided necessary implementation details in Section˜5.3.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: We follow the experimental settings of previous works to conduct experiments. Therefore, the comparison with previous works is fair.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We have provided the information of our computer resources in Section˜5.2 and Section˜A.4.3.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We ensure that our experiments comply with the NeurIPS Code of Ethics in all respects.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Please refer to our discussion on broader impacts in Section˜A.10.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks for misuse of released data or models.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We have appropriately cited related work and abide by their licenses and terms. Please refer to Section˜A.11.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: Justification: We detail the data acquisition in Appendix A.2.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: LLM is only used for writing.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Appendix
+A.1 Discussion on Motion Blur
+Image blur arises from brightness integration during exposure time and occasionally occurs in low-light conditions or high-speed motion scenarios.
+While our method does not explicitly target image blur, this does not diminish its core contributions.
+We can draw an analogy to 2D vision tasks: image deblurring addresses blur-induced degradation, while video frame interpolation tackles discontinuity from excessive inter-frame motion.
+These represent distinct challenges in computer vision.
+Similarly, in 3D reconstruction, blur-related degradation and the challenges of sparse viewpoints with pose estimation constitute separate problem domains that often co-occur in high-speed scenarios but require distinct technical solutions.
+Our work specifically addresses the latter—fundamental issues that persist regardless of blur conditions and represent critical bottlenecks in high-speed 3D reconstruction.
+Previous event-based scene reconstruction methods have primarily focused on blur reconstruction while overlooking the challenges of sparse viewpoints and inaccurate pose estimation.
+Our approach targets this gap by leveraging a novel fusion strategy to enhance traditional image-based 3DGS reconstruction, addressing fundamental limitations that affect reconstruction quality independent of blur artifacts.
+For completeness, we conducted supplementary experiments to evaluate our method’s performance under motion blur conditions.
+We extend our method to handle motion blur by incorporating the Event Double Integration (EDI) model edi , which reconstructs sharp intensity from event data.
+Specifically, we reformulate the intensity term $I_{i,0}$ in Eq.
+(5) using the EDI formulation:
+$\hat{I}_{i,0}=\frac{(2n+1)\cdot B_{i}}{\sum_{k=-n}^{n}\exp\left(C\cdot\sum_{z=0}^{k}E_{i,z}\right)},$ (19)
+where $B_{i}$ represents the blurred intensity, $E_{i,z}$ denotes the accumulated events, $C$ is the contrast threshold, and $n$ defines the temporal integration window of blur averaging.
+Experimental Setup: Experiments were performed on multiple scenes from the Tanks and Temples tnt with a frame rate of 2FPS.
+Motion-blurred frames were synthesized adopting the blur generation protocol from the GoPro-Blur dataset through gamma correction and multi-frame averaging operations.
+We average every 30 frames to simulate blurring.
+We selected EvDeblurNeRF evdeblurnerf , currently the best-performing open-source event-based deblur scene reconstruction method, as our comparison.
+Results: As demonstrated in Tab.
+5, our method achieves comparable reconstruction quality to EvDeblurNeRF while simultaneously performing pose estimation.
+Visual comparisons in Fig.
+7 show that our method renders sharper novel views under motion blur conditions.
+Note that EvDeblurNeRF requires pre-computed COLMAP poses.
+This limitation significantly restricts practical applicability in high-speed scenarios where accurate pose estimation is challenging.
+These results demonstrate that our method maintains robust performance under blur conditions while addressing the fundamental pose estimation and sparse-view challenge.
+Table 5: Comparison of different methods on pose estimation and rendering quality.
+[…]
+EvDeblurNeRF CF-3DGS 21.17 0.69 0.105 0.897
+EvDeblurNeRF COLMAP 22.58 0.70 - -
+EDI + EF-3DGS (Ours) EDI + EF-3DGS (Ours) 23.18 0.72 0.062 0.642
+Figure 7: Qualitative comparison on motion deblur scene reconstruction. Our EDI + EF-3DGS method produces sharper results comparable to EvDeblurNeRF
+A.2 Dataset
+For the synthetic dataset, following LocalRF localrf , we choose nine static scenes from the Tanks and Temples dataset, which cover both indoor and outdoor environments.
+To construct a real-world dataset, we utilized a handheld DAVIS346 event camera to capture a series of extended video sequences, simulating free camera trajectories.
+Particularly for the scenes named building, we deliberately introduced significant camera motion during acquisition to emulate realistic camera shake scenarios.
+Details about these sequences are illustrated in Table 6, where Mean.
+rotation represents the mean relative rotation angle between two adjacent frames and Max.
+rotation denotes the maximum relative rotation angle between two frames in a sequence.
+We select a 50-second segment from each sequence to highlight scenarios with free-trajectory camera movements.
+Table 6: Details of selected sequences on Tanks and Temples tnt and RealEv-DAVIS.
+Scenes Seq. Length Mean. rotation (deg) Max. rotation (deg)
+Tanks and Temples
+Auditorium 300 2.91 54.55
+Ballroom 300 5.73 179.81
+Caterpillar 300 3.18 102.37
+Church 300 2.44 37.22
+Courtroom 300 8.40 177.96
+M60 300 5.29 179.99
+Museum 300 6.05 176.98
+Panther 300 4.33 124.63
+Train 300 4.80 108.18
+RealEv-DAVIS
+building 500 5.65 67.65
+hall 500 3.88 106.73
+corner 500 3.04 96.11
+outdoor 500 2.36 77.18
+A.3 Overall Training Pipeline
+We detail the comprehensive training pipeline in Algorithm 1.
+Note that for notational simplicity, we consolidate the subscripts $(i,j)$ into a single index $k$.
+Our approach incorporates the dynamic radiance field allocation strategy from LocalRF localrf , which assigns a new radiance field when the current pose exceeds a distance threshold of 1 from the existing field.
+However, in practical implementations, we encountered the OOM(out of memory) problem with this approach.
+To address this, we adopted a modified strategy of allocating a new 3DGS every 50 frames, based on the sequence length.
+This adjustment ensures efficient memory utilization while maintaining the benefits of dynamic allocation.
+For pose optimization and scene reconstruction, we adopt strategies from CF-3DGS.
+And we incorporate event-driven brightness and motion coherence constraints, enabling robust reconstruction in challenging high-speed scenarios.
+Algorithm 1 Overall Training Pipeline.
+1:Input: event frames $\{I_{k}\}^{K}_{k=0}(k=i\cdot N+j)$.
+2:Output: camera poses $\{T_{k}\}^{K}_{k=1}$, 3DGS $\{G_{\theta_{n}}\}$.
+3:Initialize current center frame of 3DGS, $l=0$;
+[…]
+5:  $G_{\theta_{n}}$ $\leftarrow$ Initialize($I_{l}$) $\triangleright$ Following cf3dgs , initialize global 3DGS using the current center frame $I_{l}$
+6:  for k in $[l+1,min(l+50,K)]$ do $\triangleright$ Optimize the 3DGS and the poses of next 50 frames
+7:
+8:   # Pose Estimation Step
+9:   $G_{local}$ $\leftarrow$ Initialize($I_{k-1}$) $\triangleright$ Following cf3dgs , initialize local 3DGS using the previous frame $I_{k-1}$
+10:   $T_{k}$ $\leftarrow$ Optimize($\mathcal{L}_{event}$, $T_{k}$) $\triangleright$ estimate camera pose $T_{k}$ by minimizing $\mathcal{L}_{event}$ at $T_{k}$
+11:  end for
+12:
+13:  # Scene Reconstruction Step
+14:  $\{G_{\theta_{n}}\}$ $\leftarrow$ Optimize($\mathcal{L}_{event}$, $T_{[l,l+50]}$) $\triangleright$ Optimizing 3DGS $G_{\theta_{n}}$ by minimizing $\mathcal{L}_{event}$ at $T_{[l,l+50]}$
+15:  SHs of $\{G_{\theta_{n}}\}$ $\leftarrow$ Optimize($\mathcal{L}_{color}$, $T_{[l,l+50]}$)   $\triangleright$ Fixed-GS stage: optimizing SHs of 3DGS $G_{\theta_{n}}$ by minimizing $\mathcal{L}_{color}$ at RGB frames
+16:  $l\leftarrow k$ $\triangleright$ Slide the current center frame to the next center frame
+17:end while
+Table 7: Ablation study results of parameter $r$ in CMax framework
+   $r$ NVS Pose
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ $\text{RPE}_{t}\downarrow$ $\text{RPE}_{r}\downarrow$ ATE $\downarrow$
+1 22.94 0.70 0.39 0.0591 0.6829 0.4851
+3 23.96 0.72 0.36 0.0487 0.6259 0.3753
+5 23.49 0.72 0.37 0.0524 0.6431 0.4338
+A.4 Additional Experiments
+A.4.1 Detailed Experiment Results on RealEv-DAVIS
+Due to space constraints in the main text, we provide a detailed table of our method’s performance on the RealEv-DAVIS dataset in the supplementary materials, as shown in Tab. 8 and Tab. 9.
+Table 8: Quantitative Evaluations on RealEv-DAVIS.
+We select the top-performing methods from the previous evaluation.
+Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol.
+The best results are highlighted in bold.
+Methods Pose-Free Input SLOW FAST
+hall building corner outdoor hall building corner outdoor
+[…]
+[…]
+[…]
+EvDeblurNeRF $\times$ E+F 20.24 0.5762 0.55 19.35 0.5649 0.50 20.47 0.5757 0.44 22.36 0.7087 0.38 18.23 0.5043 0.65 16.41 0.4624 0.63 17.82 0.4959 0.59 19.44 0.6448 0.42
+ENeRF $\times$ E+F 21.89 0.5933 0.52 21.94 0.5918 0.36 22.01 0.5918 0.42 23.91 0.7266 0.30 20.09 0.5261 0.62 18.99 0.4903 0.50 19.61 0.5173 0.52 20.00 0.6494 0.42
+Event-3DGS $\times$ E+F 23.01 0.6085 0.47 22.45 0.6191 0.34 23.11 0.6148 0.41 25.14 0.7399 0.23 20.35 0.5328 0.62 18.52 0.4907 0.51 20.05 0.5231 0.51 21.24 0.6596 0.40
+[…]
+[…]
+Table 9: Pose accuracy on RealEv-DAVIS.
+The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100.
+Those methods which require precomputed poses are excluded.
+Methods Input SLOW FAST
+hall building corner outdoor hall building corner outdoor
+[…]
+LocalRF F 4.4017 2.5067 0.2705 4.6186 1.9409 0.4909 1.9296 2.5183 0.4177 3.4307 1.3005 0.1984 5.6061 3.2265 0.3250 7.0611 2.8722 0.6853 3.4767 3.8950 0.5693 4.7229 1.8595 0.2798
+CF-3DGS F 2.7106 0.6035 0.1459 3.0374 3.2761 0.2142 1.9437 1.4954 0.2225 2.2747 0.8053 0.0865 3.5025 0.7639 0.2635 4.6758 4.5256 0.4640 3.4950 2.2855 0.3482 3.0551 1.1198 0.1909
+Ev-Baseline E+F 0.7063 0.4386 0.1218 3.4225 1.1717 0.1988 2.1283 1.3361 0.1565 0.8647 0.3455 0.0857 0.8898 0.7734 0.1307 5.0707 1.6997 0.3166 3.6785 2.1552 0.2121 1.1645 0.5001 0.1556
+EF-3DGS(Ours) E+F 0.5262 0.3251 0.1041 2.8008 0.4633 0.1737 1.8386 1.6168 0.1864 0.4726 0.3699 0.0749 0.5424 0.3464 0.1369 3.5848 0.5589 0.2645 2.5396 2.2157 0.2029 0.5218 0.4462 0.1136
+A.4.2 Evaluating the Influence of parameter r in LEGM
+We investigate the impact of the parameter $r$ in the Contrast Maximization (CMax) framework, which determines the number of previous event frames warped to the current sampled timestamp.
+Table 7 presents the results of this ablation study.
+The optimal value for $r$ is found to be 3, yielding the best performance across all metrics for both novel view synthesis and pose estimation.
+When $r=1$, the performance degrades significantly, likely due to insufficient temporal information for accurate motion estimation.
+Increasing $r$ to 5 leads to a performance decline, though less severe than $r=1$.
+This degradation at $r=5$ can be attributed to the violation of the local linear motion assumption, which is fundamental to the CMax framework.
+A.4.3 Computational Efficiency and Impact of Subinterval Number N
+To investigate the impact of the number of subintervals N on our method, we conduct ablation studies on the Tanks and Temples dataset under the 2 FPS setting.
+The results are shown in Table 10 and Fig.
+8.
+We test the speed on RTX2080ti.
+Note that in the t-PSNR figure in Fig 8, circular nodes represent pose-free methods, while diamond-shaped nodes indicate methods that require precomputed poses.
+The results in Table 10 reveal that increasing N improves PSNR, indicating that finer temporal resolution enhances reconstruction quality.
+However, this improvement comes at the cost of an extended training time, increasing from 1.3 hours at N=2 to 7 hours at N=6.
+Importantly, our method maintains real-time rendering performance (30+ FPS), matching CF-3DGS as both utilize the same efficient 3DGS rendering framework.
+This gives our approach a significant advantage over slower NeRF-based methods.
+Even at N=2, our approach outperforms baselines in both PSNR and rendering speed.
+Event-3DGS achieves a good balance of efficiency and performance, but it requires precomputed poses.
+The setting of N=3 offers a good balance, achieving higher PSNR than CF-3DGS with comparable training time.
+However, it’s important to note that the optimal N may vary for different frame rates and motion characteristics.
+Table 10: Effect of the number of event subintervals N.
+[…]
+$N=2$ 23.56 30+ 1.3h
+$N=3$ 24.43 30+ 3h
+$N=6$ 24.81 30+ 7h
+[…]
+CF-3DGS 22.08 30+ 3h
+Figure 8: Comparison of training time vs PSNR for various methods
+Figure 9: Study of the different evalution metrics with respect to $\lambda_{cm}$
+Figure 10: Study of the different evalution metrics with respect to $\lambda_{grad}$
+Figure 11: Study of the different evalution metrics with respect to $\lambda_{PBA}$
+Figure 12: Qualitative Results of the effectiveness of $\mathcal{L}_{EGM}$.
+Figure 13: Qualitative Results of the effectiveness of $\mathcal{L}_{LEGM}$.
+Figure 14: Qualitative Results of the effectiveness of $\mathcal{L}_{PBA}$.
+A.4.4 Analysis of Loss Coefficients
+Different components of our method play crucial roles in overall performance.
+To comprehensively understand these effects and determine optimal weight settings, we conducted a detailed analysis of the loss weights: $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$.
+Contrast Maximization Coefficient $\lambda_{cm}$
+As shown in Fig.
+11, the contrast maximization coefficient $\lambda_{cm}$ significantly affects both NVS quality and pose estimation.
+We observe that a value of 0.1 achieves the best overall performance.
+Lower values (0.01) lead to decreased performance, likely due to insufficient utilization of motion information from events.
+Higher values (0.5) also result in performance degradation, possibly due to over-reliance on the contrast maximization term at the expense of other constraints.
+The optimal $\lambda_{cm}$ suggests that while event data is crucial, it should not dominate the reconstruction process.
+This balance allows our method to leverage the high temporal resolution of events without sacrificing the global supervision provided by frame data.
+Gradient-based Loss Coefficient $\lambda_{grad}$
+As shown in Fig.
+11, the gradient-based loss coefficient $\lambda_{grad}$ shows optimal performance at 0.2.
+Lower values (0.1) slightly decrease performance, while higher values (0.3, 0.5) lead to more significant drops in both NVS quality and pose accuracy.
+This is likely because $\mathcal{L}_{EGM}$ and 3DGS primarily focus on rendering absolute brightness, and excessive gradient constraints may interfere with this process.
+Therefore, a moderate $\lambda_{grad}$ value is crucial to balance gradient information with the primary rendering objectives.
+Photometric Bundle Adjustment Coefficient $\lambda_{PBA}$
+As illustrated in Figure 11, our method demonstrates remarkable stability across $\lambda_{PBA}$ values ranging from 0.1 to 0.5.
+However, while $\mathcal{L}_{PBA}$ has minimal effect on PSNR and ATE metrics, its absence leads to noticeable color distortion artifacts in visual results.
+This underscores the importance of $\lambda_{PBA}$ in maintaining visual fidelity.
+Conversely, setting $\lambda_{PBA}$ to 1.0 results in significant performance degradation, indicating that overemphasis on cross-view consistency can be counterproductive.
+A.5 Visualization of Ablation Studies
+Due to space constraints, we provide comprehensive visualizations of our ablation studies in this section.
+Our visualizations demonstrate that while CF-3DGS can produce high-quality renderings at training views, it struggles with adjacent novel views.
+The rendered depth maps expose considerable inaccuracies in scene geometry reconstruction.
+With $\mathcal{L}_{EGM}$, as shown in Fig.
+14, we observe marked improvements as it effectively harnesses continuous brightness change data captured by event cameras, compensating for inter-frame information loss.
+For $\mathcal{L}_{LEGM}$, as shown in Fig.
+14, our visualizations show that without this component, the generated Image of Warped Events (IWE) exhibits blurring effects, indicating inaccurate pose estimation.
+This is particularly evident in self-similar regions with repetitive textures, which pose additional challenges.
+$\mathcal{L}_{LEGM}$ optimizes IWE sharpness, effectively improving both rendering quality and pose estimation.
+Regarding $\mathcal{L}_{PBA}$ and Fixed-GS, as shown in Fig.
+14, our visualizations reveal that without these components, the reconstructed scene loses almost all color information.
+This occurs because event-guided optimization, with substantially more event frames than image frames, dominates scene reconstruction and suppresses color information.
+$\mathcal{L}_{PBA}$ alone, which reprojects neighboring video frames onto event frames, only partially addresses this imbalance.
+The Fixed-GS strategy separates color and structure optimization, and when combined with PBA, they complement each other to effectively resolve color distortion issues.
+A.6 Additional Challenging Scenario Experiments
+To further demonstrate event camera advantages, we constructed a FASTER scenario by selecting every 10th frame from the SLOW dataset and directly applied COLMAP pose estimation on these challenging sequences.
+We added COLMAP+3DGS as a comparison baseline.
+Notably, in the corner and outdoor scenes, several frames exhibit registration failures where the estimated poses greatly deviate from the overall video trajectory, indicating COLMAP’s limitations in challenging real-world scenarios.
+Table 11: Rendering quality comparison across different speed scenarios on RealEv-DAVIS dataset.
+Methods SLOW FAST FASTER
+[…]
+CF-3DGS 22.68 0.629 17.59 0.520 15.61 0.504
+COLMAP+3DGS 23.27 0.633 19.79 0.548 17.25 0.521
+EF-3DGS (Ours) 23.65 0.647 21.12 0.562 20.18 0.532
+As shown in Table 11, the results demonstrate that: (1) COLMAP+3DGS rendering quality degrades significantly as motion speed increase confirming pose estimation inaccuracies in challenging conditions.
+(2) Compared to CF-3DGS, our method maintains substantially better rendering quality in high-speed scenarios, highlighting our method robustness and demonstrating the potential of event cameras in high-speed scenarios.
+A.7 Necessity of Two-Stage Optimization
+To validate the necessity of our two-stage Fixed-GS training strategy, we explore an alternative approach: incorporating a reference view rendering loss $\mathcal{L}_{\text{ref}}$ at each training iteration to address the color distortion challenge without requiring a separate optimization stage.
+Given that events have much higher temporal resolution than RGB frames, the sparse color constraints from RGB images are largely overwhelmed by the abundant grayscale constraints from event data, ultimately resulting in color distortion.
+By incorporating $\mathcal{L}_{\text{ref}}$, the backpropagated gradients at each iteration would be augmented with color information from the reference view, potentially alleviating the color distortion issue.
+We conducted additional experiments comparing this approach with our two-stage strategy.
+As shown in Table 12, the results show that incorporating $\mathcal{L}_{\text{ref}}$ achieves comparable rendering quality.
+Although we cannot include visualizations here, our qualitative results confirm that the color distortion issue is effectively alleviated with this approach.
+However, this comes at a significant computational cost, increasing training time (3.8h vs 2.5h) due to additional reference view rendering at each iteration.
+Considering the modest performance difference relative to the substantial training overhead, we believe the two-stage optimization strategy offers a more efficient and practical solution.
+Table 12: Comparison of different color correction strategies.
+[…]
+Baseline 23.09 0.70 0.38 2.7h
+$\mathcal{L}_{event}+\mathcal{L}_{color}$ 24.11 0.73 0.36 3.8h
+$\mathcal{L}_{event}$ + two-stage (Fixed-GS) strategy 23.96 0.72 0.36 2.5h
+A.8 Additional Visualization
+We present additional qualitative results on both Tanks and Temples and RealEv-DAVIS.
+A.9 Limitations
+Our method combines event streams with conventional video frames, which require the input data to be time-ordered.
+So it can not handle the unordered image input.
+And, our method relies on several key parameters (e.g., $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$) that may require manual tuning for optimal performance across different speed scenarios.
+This dependency on scene-specific parameter settings could limit the method’s adaptability to diverse environments.
+Future work should explore self-adaptive parameter adjustment strategies to enhance the method’s versatility and ease of use across various reconstruction tasks.
+A.10 Broader Impacts
+To the best of our knowledge, the proposed method will not have significant negative social impact.
+The proposed scene reconstruction method can be used to reconstruct and render some wild scenes.
+Users can use the video shot by their mobile phones as input to obtain an explicit 3D asset represented by a 3D Gaussian.
+This 3D asset can be used for subsequent editing, development, secondary creation for entertainment.
+A.11 Data Availability
+The datasets that support the findings of this study are available in the following repositories: Tanks and Temples tnt at https://www.tanksandtemples.org/ under CC BY 4.0 license, The code of choosen baseline, Nope-NeRF nopenerf is available at https://github.com/WU-CVGL/BAD-Gaussians under Apache-2.0 license, LocalRF localrf is available at https://localrf.github.io/ under Apache-2.0 license, Event-3DGS event3dgs is available at https://github.com/lanpokn/Event-3DGS under Apache-2.0 license, EvDeblurNeRF evdeblurnerf is available at https://github.com/uzh-rpg/EvDeblurNeRF under Apache-2.0 license, ENeRF enerf is available at https://github.com/knelk/enerf under MIT license, CF-3DGS cf3dgs is available at https://github.com/NVlabs/CF-3DGS under Apache-2.0 license, .
+Figure 15: Qualitative comparison for novel view synthesis on Tanks and Temples. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
+Figure 16: Qualitative comparison for novel view synthesis on RealEv-DAVIS. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
+Figure 17: Pose Estimation Comparison on Tanks and Temples.
+We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (bottom colour bar) of each method.
+We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
+Figure 18: Pose Estimation Comparison on Tanks and Temples.
+We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (bottom colour bar) of each method.
+We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
+Figure 19: Pose Estimation Comparison on Tanks and Temples and RealEv-DAVIS.
+We visualise the trajectory (3D plot) and relative rotation errors $\textrm{RPE}_{r}$ (bottom colour bar) of each method.
+We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.